### v3

# Explaining Time Series Classifiers with PHAR: Rule Extraction and Fusion from Post-hoc Attributions

Abstract
Explaining machine learning (ML) models for time series (TS) classification remains challenging due to the difficulty of interpreting raw time series and the high dimensionality of the input space. We introduce PHARPost-hoc Attribution Rules—a unified framework that transforms numeric feature attributions from post-hoc, instance-wise explainers (e.g., LIME, SHAP) into structured, human-readable rules. These rules define human-readable intervals that indicate where and when decision-relevant segments occur and can enhance model transparency by localizing threshold-based conditions on the raw series. PHAR performs comparably to native rule-based methods, such as Anchor, while scaling more efficiently to long TS sequences and achieving broader instance coverage. A dedicated rule fusion step consolidates rule sets using strategies like weighted selection and lasso-based refinement, balancing key quality metrics: coverage, confidence, and simplicity. This fusion ensures each instance receives a concise and unambiguous rule, improving both explanation fidelity and consistency. We further introduce visualization techniques to illustrate specificity-generalization trade-offs in the derived rules. PHAR resolves conflicting and overlapping explanations—a common effect of the Rashomon phenomenon—into coherent, domain-adaptable insights. Comprehensive experiments on UCR/UEA Time Series Classification Archive demonstrate that PHAR may improve interpretability, decision transparency, and practical applicability for TS classification tasks by providing concise, human-readable rules aligned with model predictions.
keywords:
Explainable AI (XAI), Time Series Classification, SHAP, LIME, Anchor, Rule-Based Explanations, Post-Hoc Explainability, Expert Knowledge Extraction, Feature Attribution, Rule Fusion, Human Interpretability, Deep Learning (DL)
1 Introduction
Explaining time series classifiers presents unique challenges compared to models for images or tabular data. Time series often lack obvious semantic features. Without domain expertise, even simple attribution outputs can be difficult to interpret, which complicates the use of standard Explainable AI (XAI) tools [37]. Temporal dependencies, phase shifts, and high dimensionality further worsens explanations. When non-specialized methods are applied, results can become unstable or overly generic [12]. These issues reveal a key gap in current XAI research. Unlike non-temporal models built on intuitive, semantically rich features, time series classifiers demand explanations that both capture temporal dynamics and deliver meaningful, context-aware insights tailored to sequential data. Moreover, deep learning (DL) is widely used for time series (TS) classification, particularly in medicine and industry, because of its ability to model complex temporal patterns for tasks such as disease diagnosis and predictive maintenance. Despite strong predictive performance, these models remain hard to interpret in practice.
Aiming to address XAI challenges for state-of-the-art models such as DL, common approaches include numeric attribution methods like Permutation Importance [3], SHAP [17], and LIME [28], as well as gradient-based techniques like Saliency Maps [33] and Layer-wise Relevance Propagation (LRP) [5], all of which quantify feature contributions and trace relevance through the model’s computations. These approaches facilitate feature ranking, sensitivity analysis, and model debugging by producing developer-oriented outputs like importance scores or heatmaps. However, studies show that such numeric explanations often highlight broad regions of the input without delivering clear, context-specific insights [4, 15]. In sequential data, this tendency gives rise to "ambiguous rules", where multiple time steps or variables appear equally influential and cannot be easily distinguished. Consequently, these explanations are difficult for non-experts to interpret [14, 27], and the challenge is amplified in TS classifiers, where temporal dependencies and the high dimensionality of multivariate sequences further obscure meaningful patterns [30, 37].
We preserve the strengths of numeric-based attribution methods: precise feature ranking, post-hoc interpretability, and local fidelity, while extending them to the time series classification domain. As a first step toward addressing these challenges, rule-based approaches provide a compact, human-readable alternative to numeric attributions. We employ semi-factual reasoning to identify minimal temporal segments—interpretable intervals where perturbing feature values would not change the model’s prediction—thus generating concise, human-readable rules. Anchor rule-based explanations have demonstrated improved interpretability by extracting succinct "IF–THEN" conditions that align closely with model behavior [29]. However, such methods are seldom applied to TS data. Furthermore, rule-based methods become computationally prohibitive as sequence length increases [1]. Exhaustive search for stable "IF–THEN" conditions could be computationally infeasible, leading to a combinatorial explosion of candidate rules and temporal instability. Beyond computational infeasibility, Cognitive Load Theory, summarized by [34], emphasizes reducing extraneous demands on human working memory. Analogously, simpler and unambiguous decision rules referencing a limited number of time steps minimize cognitive effort for experts, supporting interpretability and visualization by exposing time-localized, threshold-based conditions that can be inspected on the original signal.
Our motivation stems from the persistent difficulty of interpreting high-dimensional time series models with post-hoc methods, both attribution scores and rule-based. Moreover, existing work typically assesses single explainer separately, neglecting the advantages of fusing multiple sources and systematically evaluating their quality in terms of coverage, confidence, and simplicity. Finally, the opportunity to enhance transparency by overlaying explanations on time series plots has been largely overlooked. These gaps motivate our framework, which transforms instance-level attributions into concise, human-readable rules. Rather than relying on global XAI methods based on clustering or surrogate models [21], our approach leverages numeric attributions such as SHAP [17] and LIME [28]. We adopt Anchor [29] as our baseline for comparison. Our framework unifies insights from diverse explainers, measures rule quality, and visualizes explanations directly on time series to facilitate transparent decision-making. Our framework primarily targets TS but can extend to structured data like tabular datasets.
Figure 1: Rules derived from post-hoc model explanations like SHAP.
PHAR aims to improve trust and transparency in TS classification with deep learning (DL) through three key contributions: (1) a pipeline that converts numeric post-hoc attributions into expert-level rule representations for TS; (2) a fusion mechanism that unifies rules from different explainers for better coherence and quality; (3) visualization of rules, as outlined in Figure 1. Each input time series is represented as a $T\times C$ matrix—$T$ time steps $t_{1},\dots,t_{T}$ by $C$ channels $c_{1},\dots,c_{C}$—and is processed by a deep learning classifier. Post-hoc explainers such as SHAP and LIME compute numerical attributions. In the first stage, these outputs are converted into interval rules for each instance. Rule-based explanations such as Anchor can be used directly. In the second stage, rule sets derived from different explainers are fused using methods such as union, weighted selection, and lasso-based refinement to yield a single, coherent rule per instance. Finally, each rule is visualized as vertical interval markers overlaid on the original time-series plot, enabling domain experts to inspect and validate the model’s decision logic. Crucially, the visualization conveys information beyond the textual rule: it shows rule–signal alignment in time, overlap across channels and explainers, which supports faster error checking and expert validation.
The remainder of this paper is structured as follows: Section 2 reviews related work and explains our motivation. Section 3 details our framework, covering numeric-to-rule transformation (Section 3.3) and fusion explainers (Section 3.4). Section 4 evaluates coverage, confidence, and other metrics. Section 5 summarizes findings and future directions. The Appendix A provides supplementary tables, figures, and extended results, while "Supplementary Materials" (Section B) contains statical results for Critical Difference Diagrams.
2 Related work
In this section, we review the state of the art in time series explainability that informs our PHAR framework. TS explanation methods often use surrogate models like decision trees for rule-based intervals. LIME [28], SHAP [17], and saliency maps offer numerical insights but fail to capture temporal dependencies in TS. DTW-based clustering identifies prototypical patterns, providing structural insights. Studies explore fuzzy rules, decision tree splits, and statistical impact rules, emphasizing interpretability. Some methods balance rule complexity, local fidelity, and temporal coherence, while others focus on local or global views, leaving gaps in integrating both effectively. We first examine rule-based explanation methods tailored to sequential data, and we then survey strategies for fusing multiple explainers.
Explaining TS classifiers is challenging, especially for complex models.  [21] proposed a post-hoc method using Parameterized Event Primitives (PEPs) to explain TS models. Their approach detects key events (trends, extrema), clusters them, and derives rules via decision trees. These rules capture sequential patterns, supporting fidelity to model predictions and human interpretability, while acknowledging that such rules are proxies of model behavior rather than causal explanations. This method extracts meaningful patterns while maintaining temporal dependencies and lays the groundwork for hybrid explanations that address temporal dependencies.
[32] propose a TS explanation framework combining saliency maps, instance-based examples, and rule-based methods. Saliency maps highlight key regions, instance-based methods provide prototypes and counterfactuals, and rule-based methods define logical conditions on subsequences. Integrating these methods bridges local and global explanations. This hybrid approach can improve interpretability and temporal coherence of the resulting explanations by constraining rules to compact time intervals and simple inequalities on the raw series.
Decision rules enhance interpretability in TS forecasting, especially in finance.  [4] review decision trees, fuzzy trees, and Mamdani-type rules.  [31] introduced FDT-FTS, combining fuzzy logic and decision trees for multivariate TS forecasting. Mamdani-type rules use fuzzy logic and linguistic variables for human-readable explanations.  [40] integrated them into the Neural Fuzzy Hammerstein-Wiener network. These methods balance predictive accuracy with human-readable simplicity of rules (few conditions, short intervals), supporting TS analysis. Next, we discuss decision rules for progressive decision-making.
The Belief–Rule–Based (BRB) framework [23] overcomes key weaknesses of post-hoc methods such as SHAP [17], LIME [28], and LINDA-BN [18] by structuring explanations into human-readable rules. In the Knowledge Graph Module (KGM), expert input is formalized into a causal graph that captures both temporal order and inter-feature dependencies, addressing SHAP’s and LIME’s tendency to treat each timestep or feature independently and thus lose important sequence information. The Knowledge Propagation Module (KPM) then propagates belief distributions through this graph to extrapolate feature trajectories, inherently managing missing data and uncertainty without relying on expensive local sampling that can render SHAP and LIME at times unstable. Finally, the Feature Evaluation Module (FEM) synthesizes these propagated beliefs into concise "IF–THEN" rules that highlight the most influential intervals, removing the need to interpret raw attribution weights or fuzzy inference states as in LINDA-BN and enabling straightforward expert validation.
LIMREF [25] provides local, model-agnostic explanations for global TS forecasting models. It selects similar TS via DTW and generates synthetic instances via bootstrapping. Predictions guide impact rule derivation using [39]. Rules fall into four types, including counterfactual and supportive, aiding forecast adjustments. By focusing on local TS neighborhoods, LIMREF ensures rules represent global model behavior. Its focus on local fidelity complements global and progressive methods.
[38] apply interpretable association rule mining to TS classification in maritime settings. They evaluate surrogate models for rule-based explanations of a neural network classifier. Each method generates decision rules, balancing fidelity and complexity. They analyze how support thresholds and discretization bins affect explanation quality. The authors highlights trade-offs between rule complexity, accuracy, and interpretability, showing the role of hyperparameter tuning in domain-specific rule optimization.
[20] introduced RuleXAI [^1] https://github.com/adaa-polsl/RuleXAI, an open-source rule-based explainability library for machine learning tasks. It supports classification, regression, and survival analysis by generating elementary rule conditions — simple threshold-based splits on individual features. RuleXAI provides both global and local explanations by directly learning interpretable models from data, without relying on black-box feature attribution methods. It transforms input data into binary feature conditions, reducing preprocessing complexity and ensuring high transparency. While not designed for time series data, RuleXAI’s symbolic rule extraction aligns with decision table logic and rule-based decision systems, facilitating straightforward human interpretability.
[13] explore combining numeric feature-based XAI methods in an ensemble framework, though not for TS. They propose NormEnsembleXAI, which normalizes and aggregates explanations (e.g., feature-importance, local attributions) for consistent interpretability across models. Integrating LIME, SHAP, and Integrated Gradients, they show that uniform normalization reduces attribution variability and reveals coherent explanation patterns across models.
TSProto by [7] offers a complementary approach for explaining time series classifiers by converting feature attributions (e.g., SHAP or LIME) into high-level prototypes, which are then integrated into interpretable decision trees. However, it relies on a single XAI method to identify important regions and builds rules using a separate model based on SHAP-extracted features. As a result, it remains susceptible to the Rashomon effect, since in TSProto SHAP only indicates potentially relevant regions, not how the model actually uses them in decision-making.
3 Method: Post-hoc Attribution Rules for time series
3.1 Overview of the PHAR Framework
The PHAR (Post-hoc Attribution Rules for time series) framework addresses the demand for semantically meaningful, stable, and cognitively efficient explanations in time series classification. In PHAR, we transform post-hoc, instance-level numeric feature attributions into concise, human-readable rule-based explanations and integrate outputs from multiple explainers enabling assessment of the quality of explanations and their visualization. First, PHAR selects the most salient temporal intervals—and, in multivariate settings, the most informative channels—so that rules focus on key patterns. Next, these intervals are translated into interpretable “IF–THEN” rules that preserve the model’s predicted class. An optional fusion step then merges rules from obtained from numeric attributions, Anchor explanations [29], and potentially expert rules, optimizing for coverage, confidence, and simplicity. Fusing explanations via techniques such as intersection, union, and weighted aggregation ensures faithfulness to DL classifier predictions and usability for end-users. Finally, to reduce extraneous cognitive load, PHAR visualizes each rule with semi-factual plots that contrast actual and hypothetical behaviors, highlighting how small changes would affect the model’s decision. Our framework is structured around 4 Research Questions that guide the transformation, fusion, and visualization of rule-based explanations for time series classification.
RQ1: How can numeric attributions from LIME or SHAP be transformed into precise, interpretable rule-based explanations for TS?
RQ2: How can metrics like coverage, confidence, feature count per rule, and explained instance percentage help select a single, unambiguous rule representation and reveal explanation strengths and weaknesses?
RQ3: How can fusion methods combining rules from LIME, SHAP, and Anchor generate a single optimized rule per instance? Can this transformation also refine rules within a single explainer? Whether merging multiple perspectives into one rule per instance improves coverage, confidence, and explanation ratio without excessive complexity?
RQ4: How can visualizations support the interpretability of rule-based explanations in TS models? Rules may span multiple time steps and features (in multivariate TS). We explore whether overlaying rule-based explanations on TS plots helps illustrate their structure.
Figure 2: Proposed framework with highlighted research questions RQ1…RQ4.
Figure 2 illustrates the PHAR framework with those Research Questions. A deep classifier first assigns each input series: univariate or multivariate, to a class. Post-hoc explainers SHAP and LIME yield numerical attributions that we convert into interval rules of the form $r:\;x_{c}^{(t)}\in(l_{c}^{(t)},h_{c}^{(t)}]\;\Rightarrow\;\hat{y},\quad\text{conf}(r)=\gamma,\;\text{cov}(r)=\kappa,$ where $x_{c}^{(t)}$ denotes "channel" $c$ (in case of multivariate TS) at timestep $t$, $(l_{c}^{(t)},h_{c}^{(t)}]$ is the selected interval, and $\gamma,\kappa$ are the rule’s confidence and coverage scores. This transformation addresses RQ1 (Section 3.3). In parallel, Anchor generates rule-based explanations directly. We evaluate all candidate rules using metrics such as coverage, confidence, and feature count to address RQ2 (Section 3.3.1). Then we apply multiple fusion strategies (Intersection, Union, Weighted, Lasso, Lasso global, and Best), controlled by these metrics, to select a single, optimized rule per instance (RQ3, Section 3.4). Finally, to support RQ4, we visualize each rule as vertical intervals overlaid on the original TS, highlighting feature–timestep relationships (Section 4.7 and Appendix A.4).
3.2 Rule-based explanations
In Davis et al. [10], rules are expressed in the "IF–THEN" form augmented with a certainty factor. The premise is a conjunction of predicate evaluations $p_{j}(o_{j},a_{j})=v_{j}$ for $j=1,\dots,n$, and the conclusion $q(o^{\prime},a^{\prime})=v^{\prime}_{\mathrm{out}}$ is drawn with a strength given by multiplying the aggregated premise confidence by $\mathrm{CF}_{i}$:
$R_{i}:\;\bigl(\bigwedge_{j=1}^{n}p_{j}(o_{j},a_{j})=v_{j}\bigr)\;\xrightarrow{\times\,\mathrm{CF}_{i}}\;q(o^{\prime},a^{\prime})=v^{\prime}_{\mathrm{out}}$ (1)
Here, $p_{j}$ denotes a predicate function applied to object $o_{j}$ and attribute $a_{j}$, $v_{j}$ is its value, and the certainty factor $\mathrm{CF}_{i}$ quantifies the inference reliability. Building on this idea, we define our PHAR expert-system rule format format as in [16, 24], where each rule $R_{i}$ comprises a set of logical conditions $W_{1},\dots,W_{n}$, a conclusion $W_{\mathrm{out}}$, and an inference semaphore $S_{i}$ representing the combined confidence $c_{i}$ obtained with the method described in Section 3.3.1:
$R_{i}:\;\{W_{1},W_{2},\ldots,W_{n}\}\;\xrightarrow{S_{i}}\;W_{\mathrm{out}}$ (2)
In this schema, each $W_{j}$ is a logical condition derived from feature attributions, $W_{\mathrm{out}}$ is the predicted class or decision, and $S_{i}=c_{i}$ signals the rule’s confidence.
Our approach follows established rule classifications in knowledge-based systems [22]. PHAR rules are best described as derivation and classification rules. Derivation rules define logical conditions under which a conclusion is inferred from given premises, in our case TS feature conditions leading to model predictions. Classification rules assign instances to classes based on these conditions. From the perspective of logical inference, our rules are probabilistic deductive rules. Deductive rules infer conclusions from premises. If premises hold, the conclusion follows deductively. Given a feature set and its constraints, our rules classify instances under a probabilistic confidence threshold rather than absolute certainty. Unlike classical deduction, our rules incorporate confidence, applying deterministically within probability constraints. Furthermore, our rules belong to attributive rule languages [22]. These languages encode knowledge via attributes and values, common in decision tables, rule-based systems, and databases. They enable intuitive representation and efficient manipulation. For example, the rule if $\mathrm{temperature}(t=10)>50$ and $\mathrm{pressure}(t=11)<10$, then class = failure [$c=0.85$], where $t=11$ denotes the timestep index, provides a human-readable decision boundary that replaces raw numeric attribution scores.
Converting scores from explainers which produce raw post-hoc numerical attributions, such as SHAP and LIME, into a structured rule format makes it possible to compare and merge rules from different explainers, resolve conflicts when features are correlated, and engage Domain Experts in validating and refining the symbolic rules. Our approach employs rule fusion to aggregate rule sets produced by post-hoc explainers and additionally allows to incorporate native rule-based explainers like Anchor or even rules developed directly by Domain Experts. Inspired by ensemble methods [11], we call this process fusion because it goes beyond classical bagging or boosting. Fusion refines and unifies rules by weighting them according to coverage, precision, and simplicity, and by merging overlapping or redundant conditions. The resulting fused rules conform to the same expert-system format, allowing them to be easily evaluated, adjusted, visualized, and aligned with expert knowledge to support robust domain reasoning.
We illustrate a generic rule pattern for an arbitrary instance $i$ as follows:
$R_{i}:\quad\left\{\begin{aligned} &t_{j_{1}}c_{j_{1}}\in(\ell_{j_{1}},u_{j_{1}}]\\ &t_{j_{2}}c_{j_{2}}\in(\ell_{j_{2}},u_{j_{2}}]\\ &\;\;\vdots\\ &t_{j_{k}}c_{j_{k}}\in(\ell_{j_{k}},u_{j_{k}}]\end{aligned}\right.\;\Rightarrow\;\begin{aligned} \hat{y}&="A",\\ \mathrm{CONF}(R_{i})&=\gamma,\;\gamma\in[0,1],\\ \mathrm{COV}(R_{i})&=\kappa,\;\kappa\in[0,1].\end{aligned}$ (3)
Here, each pair $t_{j_{m}}c_{j_{m}}$ denotes “channel” $c_{j_{m}}$ (in multivariate series) at timestep $t_{j_{m}}$. The indices $j_{1},\dots,j_{k}$ are selected arbitrarily from the full set of $T$ time steps and $C$ "channels", choosing the $k+1$ pairs with the greatest impact on model predictions. Each interval $(\ell_{j_{m}},u_{j_{m}}]$ has lower bound $\ell_{j_{m}}$ and upper bound $u_{j_{m}}$. $\mathrm{CONF}(R_{i})$ is the proportion of instances satisfying the rule that are correctly classified as class $A"$, and $\mathrm{COV}(R_{i})$ is the fraction of all instances that satisfy the rule. The formal derivation of rules and their CONF and COV measures is presented in Sections 3.3 and 3.3.1, while Section 4.7 provides real-world examples of rules in this notation for selected dataset.
3.3 Conversion of numeric importance values to rules
We consider a TS dataset with $N$ samples, univariate or multivariate, and a model generating multi-class predictions. In the univariate case, each feature $f$ is indexed only by timestep, e.g. $\mathrm{t}_{i}$, whereas in the multivariate case, features are indexed by both timestep and channel, e.g. $\mathrm{t}_{i}\mathrm{c}_{v}$. Let $\{X_{n}\}$ be instances with predicted classes $y_{n}$. Predicted classes serve as ground truth, as we focus on explanation consistency with DL model, not model accuracy itself. Feature-based explainers (e.g., LIME, SHAP) assign numeric importance values to each variable per timestep. Let $e_{n,f}$ denote importance values for instance $n$ and feature $f$. To extract rules, we identify a threshold $T$ (percentile-based) defining important features. Thresholds are computed per feature (local) or across all features (global).
We first compute absolute explanation values for all training instances and features. Let $E$ be the multiset of absolute values: $E=\{|e_{n,f}|:n=1,\dots,N,\;f=1,\dots,F\},$ where $e_{n,f}$ is the explanation value for feature $f$ (e.g., $\text{t}_{i}\text{c}_{v}$ for multivariate or $\text{t}_{i}$ for univariate TS) at instance $n$. We then define a global threshold $T_{global}$ by selecting a percentile $p$: $T_{global}=\text{Percentile}(E,p).$ If we apply thresholds per feature, we instead compute a separate threshold $T_{f}$ for each feature $f$: $E_{f}=\{|e_{n,f}|:n=1,\dots,N\},\quad T_{f}=\text{Percentile}(E_{f},p).$ During inference on test instances, for each feature $f$, let $e_{n,f}$ be the explanation value for instance $n$:
$|e_{n,f}|$ $\geq T_{global}\quad\text{(if using global thresholding; $\delta=True$ )},$ (4)
$\text{or}\quad|e_{n,f}|$ $\geq T_{f}\quad\text{(if using per-feature thresholding; $\delta=False$).}$
Features exceeding their threshold are important. For important features, we generate intervals preserving the model’s prediction. Stable intervals $(l_{f},u_{f}]$ are derived via controlled perturbation of each important feature and sampling around $x_{n,f}$. Let $F_{*}$ be important features for an instance, identified after thresholding. For each $f\in F_{*}$, we take the original value $x_{n,f}$ and generate perturbed samples: $x_{n,f}^{(m)}\sim\text{Uniform}(x_{n,f}-\Delta_{f},\;x_{n,f}+\Delta_{f}),$ where $\Delta_{f}$ is a perturbation range based on training explanation statistics (std. deviation × constant $\sigma_{p}$). Joint sampling over $f\in F_{*}$ captures nonlinear feature dependencies. For each perturbed sample $(x_{n,f}^{(m)})_{f\in F_{*}}$, we calculate the model’s prediction. Let $c_{ref}(n)$ be the original predicted class of instance $n$. If $c_{ref}(n)=y_{n}$, we determine $(l_{f},u_{f}]$ per feature $f$ preserving this class. Each $f\in F_{*}$ receives a stable interval: $f:x_{f}\in(l_{f},u_{f}],$ defining a rule $R_{n}$ for instance $n$:
$R_{n}:\bigwedge_{f\in F_{*}}(x_{f}\in(l_{f},u_{f}]).$ (5)
This converts numeric importance into rule conditions with stable feature intervals.
3.3.1 Confidence and coverage of newly derived rules
After deriving the final rule $R_{n}$ for instance $n$, we evaluate its explanatory quality on the test set. Consider the test dataset $\{X_{test,k}\}_{k=1}^{N_{test}}$, where $N_{test}$ is the number of test instances. We identify the subset of test instances that satisfy the rule: $S(R_{n})=\{X_{test,k}:X_{test,k}\text{ satisfies }R_{n}\}.$ To measure coverage, we compute:
$COV(R_{n})=\frac{|S(R_{n})|}{N_{test}}.$ (6)
This ratio indicates how many test instances fall into the region defined by the rule’s intervals. A higher $COV(R_{n})$ means the rule covers a larger portion of the test set.
To measure confidence, we focus on the consistency of predictions within $S(R_{n})$. Let: $T(R_{n})=\{X_{test,k}\in S(R_{n}):y_{model}(X_{test,k})=c_{ref}(n)\},$ i.e., the subset of covered test instances for which the model’s predicted class matches $c_{ref}(n)$. The confidence is then:
$CONF(R_{n})=\frac{|T(R_{n})|}{|S(R_{n})|}.$ (7)
If no test instance satisfies the rule ($S(R_{n})=\emptyset$), then $COV(R_{n})=0$ and $CONF(R_{n})$ is undefined, meaning no explanatory power. However, at least instance $n$ should satisfy the rule. If the rule covers test instances but does not match $c_{ref}(n)$, confidence decreases.
This procedure ensures coverage and confidence reflect how well derived intervals define a decision region on test data. High coverage means broad generalization; high confidence ensures alignment with model predictions.
3.3.2 Hyperparameter optimization in rule transformation step
To ensure that our rule‐based explanations achieve a balance between fidelity and simplicity, we employ Optuna [^2] Optuna: https://optuna.org/ [2] to tune key hyperparameters of this step of our framework.
Optimized hyperparameters:
  • $p$ (threshold percentile): Percentile for feature importance thresholds. Higher values prioritize top-scoring features thus results in lower quantity of important features.
  • $\delta$ (global importance indicator): Boolean for global or per-feature thresholding. If True, a single threshold $T_{global}$ is applied; otherwise, thresholds $T_{f}$ are computed per feature.
  • $\sigma_{p}$ (perturbation scale): Perturbation scale. Higher values widen the search for stable intervals.
  • $N_{p}$ (perturbation samples count): Number of perturbed samples for interval estimation. More samples improve precision but increase computational cost.
Optuna searches over the hyperparameters $p$, $\delta$, $\sigma_{p}$ and $N_{p}$ to maximize objective on a held‐out test set. During each trial, candidate configurations are evaluated by applying uniform sampling to estimate rule stability and decision region consistency, ensuring stable decision regions in the ML model’s input space. The best configuration is selected based on its performance on the test set, ensuring that the chosen thresholds generalize beyond the training data. Tuning these parameters balances feature selectivity, interval granularity, and computational cost. Table 1 (Appendix) summarizes hyperparameter ranges used in our study.
Specifically, we define an objective function maximizing coverage and confidence while penalizing empty or complex rules during numeric explanations to rule transformation step which is then feed to Optuna framework.
The objective function used in the optimization:
Each instance $n$ starts with a complex rule from an explainer (e.g., Anchor, LIME, SHAP). Each rule defines intervals for features, e.g., timestep-channel $\text{t}_{i}\text{c}_{v}$ in multivariate TS or timestep $\text{t}_{i}$ in univariate TS, denoted $F(R_{n})$. Rules yield metrics: coverage $\operatorname{COV}(R_{n})$, confidence $\operatorname{CONF}(R_{n})$, and count of features which rule specifies $|F(R_{n})|$. The initial objective function is:
$M(n)={COV}(n)\times{CONF}(n).$ (8)
If rule for instance $n$ is empty or absent, $M(n)=0$. Otherwise, penalties are applied when certain thresholds are not met.
Then we introduce thresholds $\tau_{conf}$, $\tau_{cov}$, and $\tau_{feat}$, which are chosen to balance rule quality and interpretability. $\tau_{conf}$ ensures a minimum confidence level for the rules, $\tau_{cov}$ prevents overly narrow rules by enforcing minimum coverage, and $\tau_{feat}$ limits the number of features in a rule to improve cognitive interpretability for end-users. We set $\tau_{conf}=0.5$ because scikit-learn and many binary classifiers use 0.5 as the default probability cutoff, aligning rule applicability with positive predictions above chance level and following standard model calibration practices. The coverage threshold $\tau_{cov}=0.01$ (1% support) corresponds to the minimum support filter commonly adopted in association rule mining to remove infrequent patterns and ensure rule generality. We fix $\tau_{feat}=10.0$ to limit the number of conjuncts per rule to at most ten, a choice motivated by default max-length settings in popular rule-mining libraries and grounded in human cognitive constraints defined by Miller’s Law [19] of $7\pm 2$ chunks for short-term memory.
$M(n):=\begin{array}[]{ll}M(n)\times\frac{{CONF}(n)}{\tau_{conf}},&\text{if }0<{CONF}(n)<\tau_{conf},\\[5.0pt] M(n)\times\frac{{COV}(n)}{\tau_{cov}},&\text{if }0<{COV}(n)<\tau_{cov},\\[5.0pt] M(n)\times\frac{\tau_{feat}}{Fcount(n)},&\text{if }|F(R_{n})|>\tau_{feat}.\end{array}$ (9)
These conditions are applied sequentially. The overall objective function is the average over all $N$ instances: $\bar{M}=\frac{1}{N}\sum_{n=1}^{N}M(n).$ Maximizing $\bar{M}$ finds hyperparameters balancing COV(n){COV}(n) (not too low) and CONF(n){CONF}(n) (not too low), avoiding empty solutions, and minimizing rule complexity.
3.3.3 Resolution of rules ambiguity
Several equally valid rules can be derived for a single instance and require subsequent resolution. When transforming numeric explanations into symbolic rules, different valid rules may arise depending on training data and hyperparameters. Consider three salient timestep–channel pairs $t_{j_{1}}c_{j_{1}}$, $t_{j_{2}}c_{j_{2}}$, and $t_{j_{3}}c_{j_{3}}$. The following interval rules for an arbitrary instance $i$ might be derived:
$R_{i,1}:$ $\;\bigl\{\,t_{j_{1}}c_{j_{1}}\in(\ell_{1},u_{1}]\bigr\}\;\Rightarrow\;\hat{y}=A,\;\mathrm{CONF}(R_{i,1})=\gamma_{1},\;\mathrm{COV}(R_{i,1})=\kappa_{1},$ (10)
$R_{i,2}:$ $\;\bigl\{\,t_{j_{2}}c_{j_{2}}\in(\ell_{2},u_{2}]\bigr\}\;\Rightarrow\;\hat{y}=A,\;\mathrm{CONF}(R_{i,2})=\gamma_{2},\;\mathrm{COV}(R_{i,2})=\kappa_{2},$ (11)
$R_{i,3}:$ $\;\bigl\{\,t_{j_{3}}c_{j_{3}}\in(\ell_{3},u_{3}]\bigr\}\;\Rightarrow\;\hat{y}=A,\;\mathrm{CONF}(R_{i,3})=\gamma_{3},\;\mathrm{COV}(R_{i,3})=\kappa_{3},$ (12)
Here, each $t_{j_{m}}c_{j_{m}}$ denotes "channel" $c_{j_{m}}$ (in multivariate series) at timestep $t_{j_{m}}$, each interval $(\ell_{m},u_{m}]$ has lower bound $\ell_{m}$ and upper bound $u_{m}$, and $\gamma_{m},\kappa_{m}\in[0,1]$. Although $R_{i,1}$, $R_{i,2}$, and $R_{i,3}$ yield equivalent predictions due to temporal correlations in TS, rule ambiguity arises because explainers (e.g., SHAP, LIME) combined with hyperparameter-driven optimization select different variables. This optimization, which maximizes an objective defined in Section 3.3.2, resolves ambiguity within a single explainer by choosing, for example, $R_{i,\mathrm{shap}}$ for SHAP-based attribution. However, the same procedure on LIME may produce $R_{i,\mathrm{lime}}$. To handle these divergent explanations, in Section 3.4 we propose fusing the rule set $\mathcal{R}_{i}=\{\,R_{i,\mathrm{shap}},\;R_{i,\mathrm{lime}},\;R_{i,\mathrm{anchor}}\,\}$ and selecting a single, optimized rule per instance.
3.4 Rule fusion across explainers
In rule-based systems theory [16], combining rules is called rule aggregation. Our framework calls this rule fusion, as it consolidates multiple rules into exactly one whenever possible. Unlike standard aggregation, our method also applies to a single explainer, assigning each instance a rule refined via metric optimization (e.g., Lasso method discussed later).
Let $\mathcal{F}$ be the feature set and $M$ are various explanations methods for TS classification model (e.g. rules obained from SHAP). Features are indexed by $f$, each representing a timestep-channel pair $(t,c)$ for multivaraite TS, or timestep $(t)$ for univariate TS. Explanations methods are indexed with $m$. Each method starts with multiple rules $R_{i}^{(m)}$ from different explainers $m$. Each rule contains feature conditions: $R_{i}^{(m)}=\{(f,(l_{f}^{(m)},u_{f}^{(m)}])\mid f\in F_{i}^{(m)}\},$ where $(l_{f}^{(m)},u_{f}^{(m)}]$ is an interval for feature $f$. We propose the following rule fusion approaches:
Intersection: Only conditions present in all rules per instance are kept. For each feature $f$ that appears in all $R_{i}^{(m)}$, we take the narrowest intersection of intervals:
$R_{i}^{\cap}=\bigcap_{m}R_{i}^{(m)}.$ (13)
In practice, if a feature $f$ has intervals $(l_{f}^{(m)},u_{f}^{(m)}]$ in each rule, we take:
$l_{f}^{\cap}=\max_{m}(l_{f}^{(m)}),\quad u_{f}^{\cap}=\min_{m}(u_{f}^{(m)}).$ (14)
Union: All conditions appearing in any rule are included (instance wise). For features shared among multiple rules, we take the broadest union of intervals:
$R_{i}^{\cup}=\bigcup_{m}R_{i}^{(m)}.$ (15)
If a feature $f$ has intervals from different rules, we set:
$l_{f}^{\cup}=\min_{m}(l_{f}^{(m)}),\quad u_{f}^{\cup}=\max_{m}(u_{f}^{(m)}).$ (16)
Weighted: We define a weight $w_{m}$ for each method $m$, derived from a chosen metric (e.g., confidence). We compute a weighted presence of each feature and select only those exceeding a threshold $\tau$:
$\frac{\sum_{m}w_{m}\chi_{f}^{(m)}}{\sum_{m}w_{m}}>\tau$ (17)
where $\chi_{f}^{(m)}=1$ if feature $f$ appears in $R_{i}^{(m)}$. For the selected features, intervals are combined by union.
Lasso and Lasso global: Each condition (feature-interval pair) is encoded as a binary $Z_{n,j}$, indicating if sample $n$ meets condition $j$. We then solve:
$\min_{\beta}\|y-Z\beta\|_{2}^{2}+\lambda\|\beta\|_{1}.$ (18)
Conditions with $\beta_{j}\approx 0$ are discarded, yielding sparser rules. Lasso local, Lasso for short, fits per instance; Global lasso finds conditions across all instances at once to find globally important conditions. Multiple conditions per feature are combined by union. If no rules exist for an observation in Lasso global, imputation is used. Rules are first collected per observation from different explainers (e.g., SHAP, LIME, Anchor). If no rule is assigned to an observation, the most frequent rule in the predicted class is used. If none exist, the most frequent rule overall is assigned as fallback. In Lasso local (later referred as Lasso) and other methods, observations without rules are excluded.
Best: We select the single rule that maximizes a chosen metric (e.g., confidence):
$R_{i}^{\text{best}}=\arg\max_{R_{i}^{(m)}}(\text{metric}(R_{i}^{(m)})).$ (19)
This approach avoids merging and simply picks the best-performing rule instance wise.
Finalization of resultant rules: The confidence and coverage of newly derived rules are computed as described in Section 3.3.1 and added to each rule, quantifying how well the fused rule generalizes to the data.
4 Evaluation
4.1 Model architecture
All time series datasets were classified using a unified deep neural architecture based on stacked ConvLSTM1D layers. The input consisted of time series samples with shape $(T,C)$, where $T$ denotes the number of time steps and $C$ the number of "channels". For univariate datasets, $C=1$. To capture both local and global temporal dependencies, each sequence was segmented into $n_{\mathrm{steps}}$ temporal blocks, where $n_{\mathrm{steps}}$ was chosen as the third smallest divisor of $T$, excluding $1$ and $2$. The segment length was computed as $n_{\mathrm{length}}=T/n_{\mathrm{steps}}$, resulting in an input shape of $(n_{\mathrm{steps}},n_{\mathrm{length}},C)$.
Figure 3: Architectural overview of the ConvLSTM1D-based model used in the experiments.
The model architecture included two ConvLSTM1D layers with $64$ and $32$ filters, respectively, using a kernel size of $9$ and ReLU activation. A dropout layer with a rate of $0.5$ followed, then a flattening operation producing a fixed-length embedding. This was followed by two fully connected layers: a dense layer with $100$ ReLU units and an output layer with softmax activation. Class imbalance was handled using computed class weights. The model was trained with the Adam optimizer, categorical cross-entropy loss, for $25$ epochs with a batch size of $64$. The same architecture was used across all univariate and multivariate datasets without dataset-specific tuning. A random $75:25$ train-test split was applied with stratification by class label.
A graphical summary of this architecture is shown in Figure 3, illustrating the flow from segmented TS input through ConvLSTM layers to the final classification output.
4.2 Computational complexity of PHAR
The computational cost of the proposed PHAR framework is primarily determined by two factors: the application of post-hoc explainers to time series data and the subsequent conversion of numeric attributions into symbolic rule representations and combining rules with fusion.
To provide a clear and structured analysis of computational complexity, we divide this discussion into three dedicated subsections, each corresponding to a major component of the PHAR pipeline. First, we evaluate the computational cost of the base explainers, such as SHAP, LIME, and Anchor, which generate the initial feature attributions or rule sets. Second, we analyze the cost of the numeric-to-rule transformation, which converts numeric explanations into human-readable symbolic rule intervals via perturbation-based sampling and interval search. Finally, we assess the complexity of the rule fusion process, which consolidates multiple rule sets into a single, optimized rule per instance.
4.2.1 Computational complexity of base explainers
SHAP explanations, based on kernel SHAP, have a computational complexity of $O(M^{2})$, where $M$ is the number of input features [17]. Exact SHAP computations are exponential ($O(2^{M})$), but Kernel SHAP approximates Shapley values via weighted linear regression, which still scales quadratically with $M$  [^3] https://shap.readthedocs.io/en/latest/generated/shap.KernelExplainer.html.
LIME explanations involve sampling perturbed instances and fitting a local surrogate model (typically linear regression), resulting in complexity $O(S\cdot M)$, where $S$ is the number of samples [28]. Further details are provided in the LIME documentation  [^4] https://github.com/marcotcr/lime .
Anchor explanations involve a beam search procedure over binary feature subsets, with complexity proportional to the size of the search space and sampling budget, typically $O(B\cdot S)$ where $B$ is beam width and $S$ sampling per candidate [29], see appropriate github repository [^5] https://github.com/marcotcr/anchor.
4.2.2 Computational complexity of numeric-to-rule transformation
The PHAR pipeline performs two core steps: thresholding and interval derivation; and these are wrapped inside a hyperparameter optimization loop. PHAR repeatedly applies thresholding and interval generation to evaluate different parameter configurations, and once the best hyperparameters are selected, it executes those two steps one final time to produce the final rule set. In the following subsections, we first estimate the computational cost of hyperparameter optimization (i.e., the number of repeated runs), and then analyze the individual complexities of the thresholding and interval derivation steps.
Hyperparameter optimization guides the selection of optimal threshold and perturbation settings. The search tunes four key parameters: the percentile threshold $p\in[50,99]$, a binary flag $\emph{global}\in\{\text{true},\text{false}\}$, perturbation scale $\sigma\in[0.01,1.0]$, and number of perturbation samples $S_{p}\in[1000,10000]$ (see Table 1, Appendix). We use Optuna, a Bayesian optimization framework that adaptively explores this space and prunes unpromising trials early [2]. In the worst-case scenario: if no pruning occurs—the cost scales with the number of evaluated trials $T$, which can be as high as the total number of possible combinations in the Cartesian product of parameter ranges. However, due to pruning and adaptive sampling, the actual number of full evaluations is observed to be a small fraction of $T$. For practical use, PHAR allows bypassing this stage entirely by applying default settings, e.g., $p=90$, $\emph{global}=\text{true}$, making the optimization step optional.
Thresholding requires a linear pass over the explanation values with complexity $O(N\cdot F)$, where $N$ is the number of instances and $F$ the number of features.
Interval derivation constitutes the most computationally intensive step, with complexity scaling as $O(N\cdot F^{\prime}\cdot S_{p})$, where $F^{\prime}\ll F$ denotes the average number of important features selected after thresholding, and $S_{p}$ is the number of perturbation samples generated per instance. In practice, $S_{p}$ is user-defined (typically $\leq 10000$), and both the perturbation sampling and subsequent model evaluations are highly parallelizable. This keeps the overall runtime feasible for time series datasets of moderate size, such as those commonly found in the UCR/UEA archives, where $N$ usually ranges from several hundred to a few thousand instances, and $F$ corresponds to tens or low hundreds of input features [9, 6]. This design ensures that PHAR remains computationally tractable on standard multi-core hardware.
4.2.3 Computational complexity of rule fusion
Rule fusion in PHAR uses lightweight aggregation strategies and adds negligible computational overhead compared to the previous steps.
Intersection, Union, and Weighted methods select the final rule for each instance by aggregating multiple candidate rules originating from different explainability sources. These methods rely on precomputed metrics embedded within the candidate rules (e.g., confidence, coverage) and do not require re-evaluation of the predictive model. Their computational cost scales linearly with the number of rule sources $M$ and the number of explained instances $N$, resulting in an overall complexity of $O(N\cdot M)$. Since the number of candidate rule sources per instance is typically low (e.g., $M=2$ or $3$), this step introduces negligible overhead relative to earlier phases.
Lasso involves solving a sparse linear regression problem, scaling as $O(N\cdot F_{r}^{2})$ for each instance or globally across the dataset. However, the rule matrix is highly sparse, allowing efficient optimization via coordinate descent [36]. Overall, fusion contributes marginally to total computation time, with the primary bottleneck being the initial attribution computation.
4.3 Evaluation metrics and statistical testing
We compared the baseline explainers and fusions of two or three explainers. $\bar{M}$ was applied as described in Section 3.3.2. The metric $ER$ counts instances where the explainer generated rules in proportion to all cases. $\overline{\text{CONF}}$ and $\overline{\text{COV}}$, as in Section 3.3.2, were averaged over all instances. $\overline{F(n)}$ and $Med(F(n))$ represent the mean and median number of features in generated rules (for instances with rules). Furthermore, we defined composite metrics $\overline{\text{CONF}}\cdot\text{ER}$ and $\overline{\text{CONF}}\cdot\overline{\text{COV}}\cdot ER$, which combine confidence, coverage, and rule generation rates. To compare explainers and fusion methods statistical evaluation uses non-parametric Friedman test and post-hoc analysis with Nemenyi tests. Metrics are visualized for univariate and multivariate datasets to highlight key differences. Package scikit-posthocs [35] was used.
4.4 Setup: time series, explainers, and rule fusion
The datasets used in this study come from the UCR/UEA Time Series Classification Archive. Some multivariate datasets (e.g., Libras, PenDigits) originate from the UCI ML Repository but are used via their curated UCR/UEA [^6] UCR/UEA archive: https://www.cs.ucr.edu/~eamonn/time_series_data_2018/ and https://www.timeseriesclassification.com/. Selected datasets (e.g., PenDigits) originate from UCI: https://archive.ics.uci.edu/. versions [9, 6]. Our code is available at GitHub [^7] https://github.com/mozo64/papers/tree/main/phar/notebooks in the following notebooks: UCR-explainers-lime-shap-anchor.ipynb, UCR-num2rules-optuna.ipynb, UCR-fusions-results.ipynb.
We evaluated our methods on 43 datasets, including 81.4% univariate and 18.6% multivariate TS. These datasets cover diverse domains. For each dataset, a TensorFlow model classified TS samples, followed by explanations generated using Anchor, SHAP, and LIME at the variable (>1>1 for multivariate) and time point levels. SHAP and LIME were converted into rule-based explanations using the process in 3.3.2. A separate test set, unseen during training, was used for evaluation. Table A.1.2 summarizes datasets, including the number of features used in each dataset, explainers, fusion methods and resulting fused rule sets.
Anchor, LIME, and SHAP were primary baselines, forming the foundation for fused rule sets. Anchor was used in 58.1%, LIME in 65.1%, and SHAP in 62.8% of datasets. Fusion techniques (Lasso, Weighted) refined these explanations. Lasso, Lasso global, and weighted were applied to all 43 datasets (100.0%), while Best, Intersection, and Union were used in 62.8%. The latter methods were omitted when applied to a single baseline explainer since they would replicate its results. In contrast, Lasso and Weighted optimized rules even for individual explainers. As expected, the Baseline method appeared in all datasets (100.0%) as a reference. Fused rule sets combining two or three explainers were applied in 62.8% and 23.3% of datasets, respectively.
4.5 Evaluation of fusion methods
4.5.1 Pairwise and aggregated comparisons
Performance of baseline explainers: We evaluated Anchor and rules derived from LIME and SHAP under the baseline condition, representing their original, untransformed explanations. Unlike fusion methods that optimize rule structures, baselines reflect each explainer’s native performance. To compare them, we applied the Wilcoxon signed-rank test on matched datasets, ensuring robustness for non-normal data. Table 3 presents the statistical results, while Figure 8 visualizes the distribution of key metrics.
Comparison of fusions to Anchor baseline: We applied the Wilcoxon signed-rank test to compare fusion methods with individual explainers, considering two perspectives. Tables 411 analyze fused rule sets, such as Anchor+LIME or LIME+SHAP, against the Anchor baseline. This evaluation determines whether combining multiple explainers improves key metrics.
Comparison of fusion methods to their respective baselines: Tables 1219 compare different fusion strategies, including Lasso, Weighted, and Intersection, to individual explainers like Anchor, LIME, and SHAP.
Visualization of comparisons: Figure 9 visualizes metric distributions, contrasting single explainers (orange) with fusion methods (blue) for $\bar{M}$, $\overline{\text{CONF}}\cdot ER$, $\overline{\text{COV}}$, and $F(n)$. Each boxplot includes mean, standard deviation, and median values. Additional plots detail results contrasting single explainers baseline with fusion explainers: Figure 10 ($\bar{M}$), Figure 11 ($\overline{\text{CONF}}\cdot ER$), Figure 12 ($\overline{\text{COV}}$), and Figure 13 ($F(n)$).
Median values of metrics: Table 20 reports the summary of median values, comparing rule sets derived from different explainers and theirs fusions, such as Anchor and LIME+SHAP, regardless of fusion method. Table 21 presents median values for different fusion strategies applied to explainers and theirs combinations, e.g. Intersection, Weighted, ….
4.5.2 Ranking-based and post-hoc analysis
Critical Difference Diagrams: CD diagrams rank methods across datasets based on their average performance. A lower rank generally indicates better results, as it corresponds to higher values of the evaluated metric, except for $\bar{F(n)}$, where lower values are preferred, making higher ranks more desirable. Horizontal bars connect methods that do not differ significantly at $\alpha=0.05$, while gaps denote statistically significant differences. Figure 4 presents CD diagrams for different fusion methods and baselines, ranking them according to their effectiveness across multiple metrics. Figure 5 extends the comparison to fused rule sets, showing how different explainers contribute to overall performance when used in combination. Additional figures in the appendix (Figures 14a14c) provide detailed comparisons for $\bar{M}$ , $ER$ , and $\bar{F(n)}$ , incorporating both explainer-level analysis (including individual explainers and their combinations) and specific fusion methods applied to these rule sets.
(a) $\bar{M}$
(b) $\bar{\text{CONF}}\cdot ER$
(c) $\bar{\text{CONF}}$
(d) $\bar{\text{COV}}$
(e) $\bar{F(n)}$
(f) ERER
Figure 4: Critical Difference Diagrams for selected metrics: (a) objective function $\bar{M}$, (b) $\bar{\text{CONF}}\times ER$, (c) $\bar{\text{CONF}}$, (d) $\bar{\text{COV}}$, (e) average feature count $\bar{F(n)}$, (f) explained ratio ERER. Methods joined by horizontal lines do not differ significantly at $\alpha=0.05$.
(a) $\bar{M}$
(b) $\bar{\text{CONF}}\cdot ER$
(c) $\bar{\text{CONF}}$
(d) $\bar{\text{COV}}$
(e) $\bar{F(n)}$
(f) ERER
Figure 5: Critical Difference Diagrams for selected metrics using fusion methods: (a) objective function $\bar{M}$, (b) $\bar{\text{CONF}}\times ER$, (c) $\bar{\text{CONF}}$, (d) $\bar{\text{COV}}$, (e) average feature count $\bar{F(n)}$, (f) explained ratio ERER. Horizontal lines connect methods with no statistically significant differences ($\alpha=0.05$).
Detailed tables with Friedman and Nemenyi post-hoc comparisons are provided in "Supplementary Materials" (Section B) in Tables 22– 39.
4.5.3 Compound metrics $\bar{M}$, $\overline{\text{CONF}}\cdot ER$ and $\overline{\text{CONF}}\cdot\overline{\text{COV}}\cdot ER$
Using Wilcoxon signed-rank tests, we first evaluated the baseline performance of Anchor, LIME, and SHAP, confirming slight advantage of Anchor. Among LIME and SHAP, the latter performed better for $\overline{\text{CONF}}\cdot ER$. Next, we compared fusion methods to the Anchor baseline. For $\bar{M}$ and $\overline{\text{CONF}}\cdot\overline{\text{COV}}\cdot ER$, fusions generally showed no significant gains, indicating that combining multiple explainers without considering the fusion method does not inherently improve results. However, when each fusion strategy was compared to its respective single-explainer baseline, we observed consistent improvements across all three compound metrics, with the best results from Lasso, Weighted, and Best. Other methods yielded comparable, worse, or statistically insignificant results. Friedman tests confirmed significant differences ($p<0.05$) among both fusion methods and explainer combinations. Nemenyi post-hoc analysis further highlighted differences in performance, particularly for the best-performing methods. Critical Difference Diagrams (CDD) reinforced these findings, ranking Lasso, Weighted, and Best as the top fusion strategies. Among explainer combinations, Anchor and its fusions performed best, with LIME+SHAP achieving comparable results.
[…]
Wilcoxon signed-rank tests showed SHAP outperforming LIME in $\overline{\text{CONF}}$, while Anchor+SHAP exceeded Anchor. Most multi-explainer fusions improved $\overline{\text{COV}}$, but no significant differences were found for $ER$. Fusion methods Best, Lasso, Union, and Weighted consistently improved all three metrics over baselines. Friedman tests confirmed significant differences ($p<0.05$), and Nemenyi post-hoc analysis validated the superior performance of these fusion methods. CDD illustrate these trends, ranking Best, Lasso, Union, and Weighted highest for $\overline{\text{CONF}}$, $\overline{\text{COV}}$, and $ER$. Lasso global excelled in $\overline{\text{COV}}$. Among explainer combinations, Anchor-based fusions ranked highest, with LIME+SHAP performing well for $\overline{\text{CONF}}$ and $ER$.
4.5.5 Features Count $\overline{F(n)}$ and $\mathrm{Med}(F(n))$
Wilcoxon signed-rank tests showed Anchor using the fewest features, with LIME generating fewer than SHAP. Against the Anchor baseline, all fused explainers had higher $\overline{F(n)}$ and $\mathrm{Med}(F(n))$. Among fusion methods, Intersection and both Lasso variants minimized feature count, while Union ($42.6$ vs. $23.6$) and Best ($27.5$ vs. $23.6$) increased it. Friedman and Nemenyi tests confirmed significant differences across methods and explainer combinations. CDD ranked Lasso, Lasso global, and Intersection as the most efficient in minimizing $\overline{F(n)}$, while Anchor, SHAP, and Anchor+SHAP ranked best in that regard among explainer combinations.
4.6 Summary of evaluation results
Anchor consistently performed best in $\bar{M}$ and minimizing feature count, but was not dominant in $\overline{\text{CONF}}$, $\overline{\text{COV}}$, and $ER$. Among explainers, SHAP outperformed LIME in $\overline{\text{CONF}}$, and Anchor+SHAP surpassed Anchor. Most multi-explainer fusions improved $\overline{\text{COV}}$, while no significant differences were found for $ER$. Fusion methods had mixed effectiveness and choosing the right method is crucial. In compound metrics, Lasso, Weighted, and Best consistently improved performance. Some methods, such as Union, increased feature count significantly, while Lasso and Intersection minimized it. CD Diagrams confirmed that fusions often outperformed individual explainers. The best methods depended on the metric: Lasso, Weighted, and Best excelled in $\bar{M}$ and compound metrics, while Lasso global was particularly effective in $\overline{\text{COV}}$. Among explainer combinations, Anchor-based fusions ranked highest, but LIME+SHAP also performed well in $\overline{\text{CONF}}$ and $ER$.
4.7 Qualitative evaluation
We illustrate PHAR on the ECG200 dataset (two classes) using observation 21. Because the data are univariate, the channel index is omitted and rules refer directly to time-step indices. For LIME and SHAP we convert per–time-step attributions into contiguous interval constraints using the procedure in Sec. 3.3; Anchor provides a native rule set. We then form rule sets by three fusion operators (Sec. 3.4): union (all intervals from the inputs are pooled into one conjunctive rule), lasso (a sparse re–selection of intervals under an $\ell_{1}$ penalty), and best (a single rule is chosen by a deterministic ranking: confidence has top priority; in a tie a fixed explainer order breaks ties). [^8] Other tie–breakers, e.g. adding a sparsity term, are possible but not used here. All values are standardised $z$-scores relative to the training set. Thus thresholds can be read as deviations from the mean, which makes the intervals actionable: one can see how much the signal would need to move to break or satisfy a condition. In practice, the interval conditions ($t_{i}\in[\ell,u]$) are contextualised as characteristic segments, such as an initial dip, a mid‑sequence plateau or a terminal return to baseline—which makes the explanation more natural than a bare list of thresholds.
We recommend a two–step reading path. First, inspect the visualised intervals to understand where in the trace the decision is anchored. Second, examine the explicit rule statements to see the exact inequalities. We follow this order below.
Figure 6 shows the rule intervals for observation 21. The blue curve is the signal. Red dashed segments mark the time intervals required by each rule $R_{21}$. The numbers in each panel report confidence and coverage on the test set. Panels (a)–(c) present single-explainer benchmarks: Anchor, LIME, and SHAP. Panels (d) and (f) show union fusions: Anchor+SHAP and LIME+SHAP. Panels (e) and (h) show lasso fusions: LIME+SHAP and Anchor+LIME+SHAP. Panel (g) reports the best selection over all explainers. We next discuss each panel in turn.
(a) Anchor (benchmark). Two semi-infinite thresholds at $t_{42}$ and $t_{51}$ localise the decision in the mid/late part of the trace. The rule is extremely compact (high confidence, moderate coverage), but the one-sided nature of the thresholds provides little temporal extent, which limits shape specificity.
(b) LIME (benchmark). Many short intervals concentrate in the mid section ($\approx\!t_{36}$$t_{56}$) and late flat portion, and LIME also flags the early inflection around $t_{16}$$t_{17}$ and another change near $t_{22}$. This pattern closely tracks local shape variations and attains full confidence at low coverage.
(c) SHAP (benchmark). SHAP emphasises the initial drop ($t_{4}$$t_{6}$, $t_{10}$$t_{11}$) and the local peak/plateau around $t_{22}$$t_{30}$. It therefore complements LIME by assigning more weight to the early and early-mid morphology.
(d) Fusion (union): Anchor+SHAP. The union operator pools all conditions from both inputs into a single conjunctive rule. In this instance the union has the same confidence and coverage as SHAP because every test case that satisfies SHAP’s intervals already satisfies Anchor’s mid/late thresholds, so the added Anchor conditions are redundant and the covered set is unchanged. Since adding conjunctive conditions cannot increase coverage and does not decrease it here, confidence also remains the same. The extra Anchor intervals mainly make the mid section explicit in the panel.
(e) Fusion (lasso): LIME+SHAP. Lasso reduces cognitive load by pruning overlapping or lower-priority intervals under the $\ell_{1}$ selection criterion. The retained set remains balanced: it keeps informative early (from SHAP), mid (from both), and late segments (from LIME), yielding a concise yet representative rule.
(f) Fusion (union): LIME+SHAP. Pooling both explainers yields the most extensive set of intervals along the trace—dense mid/late coverage from LIME together with SHAP’s early emphasis. This gives a simple reading: early evidence is provided by SHAP and mid/late evidence by LIME, allowing a practitioner to probe "WHAT-IF" edits across substantial part of waveform. Adding intervals in a conjunctive rule can maintain or reduce coverage, but not inflate it.
(g) Fusion (best): Anchor+LIME+SHAP. For this observation, the best selector returns exactly the LIME (benchmark) rule. LIME achieves the highest confidence at the attained coverage. In this case SHAP attains the same metrics, but our deterministic tie-break (fixed explainer order) selects LIME, so SHAP is not chosen.
(h) Fusion (lasso): Anchor+LIME+SHAP. Lasso selects a sparse subset drawn from all three sources, without favouring a single explainer. The resulting rule captures dispersed early, mid and late segments while remaining sparse (compact).
To sum up, Anchor provides the shortest and most readable rule, but it is less useful in visual analysis: its one-sided, semi-infinite thresholds highlight isolated time indices rather than contiguous segments, so they convey little about waveform duration or shape. LIME and SHAP emphasise different phases (mid/late vs. early/early-mid). Union fusions offer the broadest spatial map of supporting intervals, whereas lasso fusions distil them to a compact subset. The best fusion method returns a single rule according to the stated priorities; in this case the selected rule is LIME, tied with SHAP on the metrics.
(a) Anchor (benchmark)
(b) LIME (benchmark)
(c) SHAP (benchmark)
(d) Fusion (union): Anchor + SHAP
(e) Fusion (lasso): LIME + SHAP
(f) Fusion (union): LIME + SHAP
(g) Fusion (best): Anchor + LIME + SHAP
(h) Fusion (lasso): Anchor + LIME + SHAP
Figure 6: Observation 21 from ECG200 dataset (univariate). Rule intervals produced by single explainers (Anchor, LIME, SHAP) and by various fusion strategies. The blue curve is the signal; red dashed bands mark the value intervals used by rule $R_{21}$ for each method. Confidence and coverage are printed within each panel. Definitions of fusion methods are provided in Sec. 3.4.
4.7.1 Counterfactuals: testing rule boundaries
Counterfactuals help assess whether an instance-level rule genuinely separates classes in ways that are meaningful to practitioners. In our setting, a counterfactual (CF) is defined operationally: it is a test observation drawn from the opposite class and overlaid as a dotted trace. This construction is diagnostic rather than causal: it provides a direct visual contrast that lets experts see where and how the rule would be violated by an instance that should be rejected.
We use CFs to clarify what the rule demands and which time segments carry the decision. All series are shown in the same normalised scale. Red dashed segments indicate the admissible value ranges at specific, constrained time indices $t$; these may be two-sided intervals or one-sided thresholds. The blue curve (target observation) satisfies all marked ranges. The grey dotted CF does not: whenever the dotted trace exits a red segment at a constrained $t$, the rule is broken and the CF remains in the opposite class. A small box on each panel reports the rule’s $\mathrm{CONF}$ and $\mathrm{COV}$, which contextualise how reliable and how frequent the rule is across the test set. This combination—contrast against a CF, explicit intervals, and summary metrics—yields a compact, expert-friendly reading of the decision boundary.
(a) Anchor (benchmark): obs. 21 vs. CF 38
(b) SHAP (benchmark): obs. 21 vs. CF 3
(c) LIME+SHAP (lasso): obs. 21 vs. CF 47
Figure 7: Examples illustrate three Counterfactuals for instance 21: (a) Anchor (benchmark) with CF instance 38; (b) SHAP (benchmark) with CF instance 3; and (c) LIME+SHAP with lasso fusion with CF instance 47.
The three panels tell a coherent story: CF leaves the admissible bands, which results in a class flip. In panel (a) Anchor (benchmark) imposes two mid-series intervals, around $t\!\approx\!43$ and $t\!\approx\!50$, with narrow bands slightly above zero; the target trace remains inside both, whereas the CF crosses them from above and thus violates the rule. In panel (b) SHAP (benchmark) concentrates on the early rebound after the initial dip ($t\!\approx\!20\text{–}27$) and requires a short positive excursion; the target exhibits this burst, while the CF stays lower and repeatedly exits the band. In panel (c) LIME+SHAP (lasso) selects a compact sequence of mid-section intervals ($t\!\approx\!39\text{–}60$) centred near zero with a slight positive offset; the target stays within these tight corridors, whereas the CF drifts outside at several constrained times.
4.7.2 Rule listings for the panels
Guided by the rule intervals in Fig. 6 and Fig. 7, we now give the corresponding rule statements for panels (a)–(h). Each listing follows the same notation and uses standardised $z$-scores. Presenting visuals before equations aids time-series interpretation: the reader can first locate the relevant phases of the waveform and then verify the precise inequalities. Showing the inequalities after the visuals keeps the focus on where the decision is anchored and then makes it precise enough to audit or run simple "WHAT-IF" perturbations. This also facilitates counterfactual checks—one can perturb the signal within a highlighted window to test whether the rule would be violated. Panel labels are retained so the reader can cross-reference each rule with its subfigure.
a) Anchor (benchmark)
$R_{21}:\;\left\{\begin{aligned} &t_{42}\leq 0.58\\ &t_{51}\leq 0.23\end{aligned}\right$.
$\hskip 2.12503pt\Rightarrow\;\begin{aligned} \hat{y}=\text{"0"},\\ \mathrm{CONF}(R_{21})=0.90,\\ \mathrm{COV}(R_{21})=0.20.\end{aligned}$
c) SHAP (benchmark)
[…]
$\hskip 2.12503pt\Rightarrow\;\begin{aligned} \hat{y}=\text{"0"},\\ \mathrm{CONF}(R_{21})=1.00,\\ \mathrm{COV}(R_{21})=0.02.\end{aligned}$
d) Union of Anchor + SHAP
[…]
$\hskip 2.12503pt\Rightarrow\;\begin{aligned} \hat{y}=\text{"0"},\\ \mathrm{CONF}(R_{21})=1.00,\\ \mathrm{COV}(R_{21})=0.02.\end{aligned}$
b) LIME (benchmark)
[…]
$\hskip 2.12503pt\Rightarrow\;\begin{aligned} \hat{y}=\text{"0"},\\ \mathrm{CONF}(R_{21})=1.00,\\ \mathrm{COV}(R_{21})=0.02.\end{aligned}$
f) Union of LIME + SHAP
[…]
$\hskip 2.12503pt\Rightarrow\;\begin{aligned} \hat{y}=\text{"0"},\\ \mathrm{CONF}(R_{21})=1.00,\\ \mathrm{COV}(R_{21})=0.02.\end{aligned}$
g) Best of Anchor + LIME + SHAP
[…]
$\hskip 2.12503pt\Rightarrow\;\begin{aligned} \hat{y}=\text{"0"},\\ \mathrm{CONF}(R_{21})=1.00,\\ \mathrm{COV}(R_{21})=0.02.\end{aligned}$
e) Lasso of LIME + SHAP
[…]
$\hskip 2.12503pt\Rightarrow\;\begin{aligned} \hat{y}=\text{"0"},\\ \mathrm{CONF}(R_{21})=1.00,\\ \mathrm{COV}(R_{21})=0.02.\end{aligned}$
h) Lasso of Anchor + LIME + SHAP
[…]
$\hskip 2.12503pt\Rightarrow\;\begin{aligned} \hat{y}=\text{"0"},\\ \mathrm{CONF}(R_{21})=1.00,\\ \mathrm{COV}(R_{21})=0.02.\end{aligned}$
4.7.3 Dataset-level visual summary (ECG200 and Plane datasets)
Beyond the single-trace analysis in Sec. 4.7, Appendix A.4 presents a dataset-level visual catalogue for ECG200 (Appendix A.4.1) and Plane (Appendix A.4.2). For each explainer (Anchor, LIME, SHAP) and their fusions (union, lasso, best), we show curated per-class exemplars (diverse instances, prototype, class mean) with the same red-window overlay used throughout. This complements the pointwise analysis by highlighting recurrent early/mid/late concentrations at the dataset level.
5 Conclusion
This work introduces a novel rule-based explainability framework tailored for time series (TS) classification tasks, addressing the current gap in interpretable AI solutions for temporal data. To the best of our knowledge, this is the first systematic study transforming post-hoc, instance-based numerical feature attributions from SHAP and LIME explainers into structured, human-readable rule representations for TS. Our approach uniquely integrates and compares these rule-based explanations with Anchor, a native rule-based explainer, across a comprehensive benchmark of 43 TS datasets from the UCR/UEA Time Series Classification Archive. We further contribute a modular rule fusion pipeline, combining multiple explainers into a single, optimized rule per instance while ensuring high fidelity and coverage. This work is accompanied by evaluation of multiple rule aggregation strategies, along with an open-source codebase and practical visualization techniques to facilitate adoption in real-world applications.
We transform numeric feature attribution explainers (LIME, SHAP) into rule-based TS explanations, integrating them with Anchor. Our post-hoc framework explains instances post-training, crucial for TS tasks like anomaly detection and decision support (RQ1). Feature-to-rule conversion balances complexity, improving coverage, confidence, and ensuring interpretability for each instance. Optimizing thresholds and perturbations refines structured, unambiguous rules, improving interpretability. Reliable rules align with model decisions, balancing confidence and coverage (RQ2).
SHAP and LIME perform comparably to Anchor, despite the latter being natively rule-based. While Anchor incurs high computational cost on long TS, converting SHAP attributions into interval rules provides a more efficient alternative that achieves competitive performance when optimized via fusion. Moreover, rules derived from our numeric-to-rules transformation use closed intervals $(\ell,u]$ for each timestamp-channel condition, whereas Anchor often produces one-sided bounds. Closed intervals are especially important in TS, as they precisely delimit both lower and upper thresholds, ensuring that temporal segments are fully specified, reducing ambiguity in sequential explanations, and allowing for clearer visualization.
Our framework consolidates knowledge from multiple explainers (RQ3). Unlike bagging or boosting, our post-hoc fusion refines single or multiple explainers into a coherent rule set, aligning with expert system principles. Fusion improves rule reliability, reduces ambiguity, and enhances decision-making. Moreover, PHAR supports sparsity optimization. We measure sparsity with the rule-length metric $F(n)$ and report both its mean and median. This is important given human working-memory limits (the 7 $\pm$ 2 principle) [19], as discussed earlier when defining the objective function. Our method enables this, when desired, by offering resolution modes that promote concise, non-overlapping rules, and by using fusion methods such as lasso and intersection. This is also illustrated by the ECG200 example discussed in our qualitative evaluation.
Rules enhance TS visualization, aiding expert interpretation (RQ4). Anchor uses fewer features but produces broad intervals, missing finer temporal details. SHAP and LIME involve more features, creating narrower, potentially more human-readable intervals that retain predictive alignment while avoiding redundant or overlapping conditions. Application-dependent, broader feature sets or refined intervals enhance explanation resolution.
Interpretability of fused rules depends on simplicity (fewer features) and uniqueness (avoiding redundancy), optimized via key metrics. Rule fusion removes redundancy, ensuring a structured, coherent set of explanations. Results show that both single-explainer (SHAP, LIME) and fusion methods achieve high confidence, coverage, and explanation ratio. Lasso, Weighted, and Best refine rules, often improving clarity and consistency, while visualization enhances the illustration of rule applicability.
5.1 Limitations
The PHAR framework effectively transforms numeric feature attributions into rule-based explanations, but several limitations remain.
  1. 1.
    Rule quality within PHAR depends on explainers such as SHAP and LIME, which rely on perturbation-based sampling and may produce unstable attributions, particularly in noisy or highly dynamic time series.
  2. 2.
    Computational complexity represents a secondary limitation; while the transformation step in PHAR involves hyperparameter tuning, we have shown in Section 4.2 that the practical cost remains acceptable due to parallelization, the optional nature of tuning, and the complexity of the fusion step, with the exception of the more computationally intensive Lasso global variant.
  3. 3.
    Our evaluation of PHAR employed a predefined set of metrics, and further exploration of alternative rule selection criteria could enhance the flexibility of rule generation.
  4. 4.
    Interpretability within PHAR may be challenged in high-dimensional time series, where resulting rules risk becoming overly specific; however, this issue is partially mitigated by the dedicated visualization techniques developed in PHAR, which follow a semifactual design and aim to present rule intervals in a clearer, more interpretable manner. Future work could further address this limitation by incorporating feature selection mechanisms or additional constraints to promote more generalizable rule representations.
  5. 5.
    While our visualization experiments within PHAR demonstrated practical applicability, we did not conduct evaluations with Domain Experts, and the relationship between rule complexity and perceived clarity remains to be systematically investigated.
  6. 6.
    Finally, although fusion in PHAR balances rule importance, it does not fully eliminate potential conflicts between different explainers. Additionally, we did not explicitly analyze rule similarity or stability, which could be particularly relevant in time series contexts where neighboring time steps may influence model predictions in a correlated manner.
5.2 Future work
While our framework effectively transforms numeric feature attributions into rule-based explanations, several directions remain open.
  1. 1.
    One of the important directions is extending PHAR to counterfactual explanation generation, by identifying and presenting the nearest time series belonging to an alternative class. This would allow Domain Experts to specify transitions from the current predicted class to any arbitrary target class (e.g., in predictive maintenance or clinical decision support), thereby providing actionable "WHAT-IF" insights.
  2. 2.
    Further research should assess rule stability across similar instances, determining whether minor input variations lead to consistent explanations.
  3. 3.
    Integrating expert feedback could refine rule usability, bridging automated rule extraction with human interpretability. One approach is incorporating symbolic reasoning systems like HeaRTDroid [8]  [^9] https://heartdroid.re/, a lightweight rule engine for expert systems, enabling a hybrid approach that merges statistical learning with structured decision-making.
  4. 4.
    For improvement of interpretability, it will be worth investigating the trade-off between feature count and interval width.
  5. 5.
    Usage of advanced visualization methods, such as interactive overlays or multi-dimensional interval representations could be beneficial. Related direction is DeepVATS [26]  [^10] https://github.com/vrodriguezf/deepvats, which integrates DL with interactive visualization of latent space projections for TS, similar to TensorFlow Embeddings Projector, aiding in anomaly detection and structure analysis.
  6. 6.
    Finally, expanding the framework beyond TS classification to areas like sequential decision-making or tabular datasets with temporal dependencies could validate its broader applicability.
[…]
Supplementary information Additional tables supporting the statistical comparison of rule-based explainers and fusion methods via Critical Difference Diagrams are provided in "Supplementary Materials" (Section B).
[…]
Acknowledgements This paper is part of a project that has received funding from the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No. 101120406. The paper reflects only the authors’ view and the European Commission is not responsible for any use that may be made of the information contained herein.
Contribution of Maciej Mozolewski for the research for this publication has been supported by a grant from the Priority Research Area (DigiWorld) under the Mark Kac Center for Complex Systems Research Strategic Programme Excellence Initiative at Jagiellonian University.
We thank all contributors of open-source libraries (e.g., TensorFlow, Optuna) and the maintainers of the UCR/UEA Time Series Classification Archive for curating and providing the datasets. We used OpenAI’s GPT to assist in writing and refining the text and generating code snippets, while all conceptual ideas and research contributions were developed by the authors.
Appendix A
A.1 Additional Tables
A.1.1 Hyperparameter for Numeric-to-Rules Transformation
Table 1: Hyperparameter Ranges for Optimization
Hyperparam. Range/Values Description
pp 50 to 99 (step=1) Percentile for determining feature importance thresholds.
$\delta$ {True, False} Global vs. per-feature threshold application.
$\sigma_{p}$ 0.01 to 1.0 (continuous) Scale of perturbation for interval estimation.
$N_{p}$ 1,000 to 10,000 (step=1,000) Number of perturbed samples for interval estimation and confidence evaluation.
  • $p$ - threshold percentile, $\delta$ - global importance indicator, $\sigma_{p}$ - perturbation scale, $N_{p}$ - perturbation samples count
A.1.2 Experimental Time Series, Explainers and Fused Rule Sets
Methods Baselines Fused rule sets
Dataset Type Acc #F Dim #B #E
Best
Union
Lasso global
Lasso
Weighted
Intersection
Anchor
SHAP
LIME
Anchor + SHAP
Anchor + LIME
LIME + SHAP
Anchor + LIME + SHAP
ArticularyWordRecognition m 96,53 1296 $144\times 9$ 2 12
AtrialFibrillation m 37,50 1280 $640\times 2$ 2 12
BasicMotions m 100,$0$ 600 $100\times 6$ 2 12
Cricket m 93,33 7182 $1197\times 6$ 2 12
ERing m 98,67 260 $65\times 4$ 2 12
Epilepsy m 91,30 618 $206\times 3$ 2 12
Libras m 63,33 90 $45\times 2$ 1 3
PenDigits m 99,24 16 $8\times 2$ 1 3
Adiac u 29,59 176 $176\times 1$ 2 12
BME u 86,67 128 $128\times 1$ 2 12
Beef u 53,33 470 $470\times 1$ 2 12
BeetleFly u 80,$0$ 512 $512\times 1$ 2 12
BirdChicken u 50,$0$ 512 $512\times 1$ 2 12
CBF u 100,$0$ 128 $128\times 1$ 2 12
Chinatown u 98,90 24 $24\times 1$ 3 33
CricketX u 61,54 300 $300\times 1$ 1 3
DiatomSizeReduction u 100,$0$ 345 $345\times 1$ 2 12
DistalPhalanxOutlineAgeGroup u 77,78 80 $80\times 1$ 1 3
DistalPhalanxOutlineCorrect u 78,08 80 $80\times 1$ 1 3
DodgerLoopDay u 55,$0$ 288 $288\times 1$ 3 33
DodgerLoopGame u 87,50 288 $288\times 1$ 3 33
ECG200 u 86,$0$ 96 $96\times 1$ 3 33
ECG5000 u 91,52 140 $140\times 1$ 1 3
ECGFiveDays u 100,$0$ 136 $136\times 1$ 3 33
FaceFour u 96,43 350 $350\times 1$ 1 3
GunPoint u 86,$0$ 150 $150\times 1$ 2 12
GunPointAgeSpan u 88,50 150 $150\times 1$ 3 33
GunPointMaleVersusFemale u 100,$0$ 150 $150\times 1$ 2 12
GunPointOldVersusYoung u 100,$0$ 150 $150\times 1$ 2 12
InsectWingbeatSound u 66,91 256 $256\times 1$ 1 3
MedicalImages u 68,53 99 $99\times 1$ 1 3
MiddlePhalanxOutlineAgeGroup u 78,42 80 $80\times 1$ 1 3
PhalangesOutlinesCorrect u 67,07 80 $80\times 1$ 1 3
Plane u 96,23 144 $144\times 1$ 3 33
PowerCons u 100,$0$ 144 $144\times 1$ 1 3
ProximalPhalanxTW u 48,68 80 $80\times 1$ 3 33
SmoothSubspace u 96,$0$ 15 $15\times 1$ 1 3
Strawberry u 77,64 235 $235\times 1$ 1 3
SyntheticControl u 89,33 60 $60\times 1$ 1 3
Trace u 68,$0$ 275 $275\times 1$ 1 3
TwoPatterns u 99,92 128 $128\times 1$ 2 12
UMD u 93,33 150 $150\times 1$ 3 33
Wafer u 99,83 152 $152\times 1$ 3 33
Table 2: Dataset summary with baselines, methods, and fusion methods. Type is represented as u (univariate) or m (multivariate). Column Acc denotes Accuracy of the ML model. Column #F denotes the total number of input features, and Dim denotes the input time series dimensionality (time $\times$ variables). Columns #B and #E indicate the number of baselines and ensembles for each dataset.
A.1.3 Wilcoxon Signed-rank Tests
Table 3: Wilcoxon signed-rank test results comparing baselines (single explainers: Anchor, LIME, SHAP) for each metric. Columns include medians for each pair, test statistic ($W$), p-value ($p$), and the number of compared pairs (NN). Statistically significant results are marked with asterisks: ($p<0.05$), ∗∗ ($p<0.01$), ∗∗∗ ($p<0.001$).
Metric Explainers NN Median1 Median2 $W$ $p$
$\bar{M}$ Anchor vs LIME 10 16.89 4.95 2.0 0.006∗∗
Anchor vs SHAP 10 16.89 6.86 4.0 0.014
LIME vs SHAP 27 3.30 2.56 177.0 0.79
$\bar{{CONF}}\cdot{ER}$ Anchor vs LIME 10 83.43 65.75 20.0 0.49
Anchor vs SHAP 10 83.43 86.99 24.0 0.77
LIME vs SHAP 27 63.70 89.17 55.0 < 0.001∗∗∗
$\bar{{CONF}}\cdot\bar{{COV}}\cdot{ER}$ Anchor vs LIME 10 17.48 6.03 5.0 0.019
Anchor vs SHAP 10 17.48 7.82 6.0 0.027
LIME vs SHAP 27 7.83 6.24 154.0 0.41
$\bar{{CONF}}$ Anchor vs LIME 10 0.88 0.88 16.0 0.28
Anchor vs SHAP 10 0.88 0.94 24.0 0.77
LIME vs SHAP 27 0.82 0.92 60.0 0.001∗∗
$\bar{{COV}}$ Anchor vs LIME 10 0.20 0.12 12.0 0.13
Anchor vs SHAP 10 0.20 0.09 11.0 0.11
LIME vs SHAP 27 0.13 0.10 113.0 0.07
ER{{ER}} Anchor vs LIME 10 98.03 88.10 17.0 0.89
Anchor vs SHAP 10 98.03 100.00 10.0 0.92
LIME vs SHAP 27 86.11 100.00 62.0 0.06
$\bar{F(n)}$ Anchor vs LIME 10 1.94 13.46 1.0 0.004∗∗
Anchor vs SHAP 10 1.94 7.12 1.0 0.004∗∗
LIME vs SHAP 27 19.98 32.04 137.0 0.22
Med(F(n)){Med}(F(n)) Anchor vs LIME 10 2.00 8.00 4.0 0.014
Anchor vs SHAP 10 2.00 6.50 1.5 0.013
LIME vs SHAP 27 9.50 21.00 89.0 0.048
Table 4: Wilcoxon signed-rank test results comparing explainers (and their fused rule sets) to Anchor (Baseline) for $\bar{M}$.
Explainers set $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Anchor+LIME vs Anchor 10 9.94 17.29 0.00 0.002∗∗
Anchor+LIME+SHAP vs Anchor 10 7.60 17.29 3.00 0.010∗∗
Anchor+SHAP vs Anchor 10 11.03 17.29 5.00 0.019
LIME vs Anchor 10 4.65 17.29 0.00 0.002∗∗
LIME+SHAP vs Anchor 10 5.79 17.29 0.00 0.002∗∗
SHAP vs Anchor 10 8.32 17.29 0.00 0.002∗∗
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 5: Wilcoxon signed-rank test results comparing explainers (and their fused rule sets) to Anchor (Baseline) for $\bar{{CONF}}\cdot{ER}$.
Explainers set $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Anchor+LIME vs Anchor 10 74.08 68.71 20.00 0.49
Anchor+LIME+SHAP vs Anchor 10 69.32 68.71 27.00 1.00
Anchor+SHAP vs Anchor 10 77.28 68.71 15.00 0.23
LIME vs Anchor 10 59.68 68.71 15.00 0.23
LIME+SHAP vs Anchor 10 73.12 68.71 26.00 0.92
SHAP vs Anchor 10 70.64 68.71 21.00 0.56
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 6: Wilcoxon signed-rank test results comparing explainers (and their fused rule sets) to Anchor (Baseline) for $\bar{{CONF}}\cdot\bar{{COV}}\cdot{ER}$.
Explainers set $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Anchor+LIME vs Anchor 10 11.36 18.64 4.00 0.014
Anchor+LIME+SHAP vs Anchor 10 9.02 18.64 4.00 0.014
Anchor+SHAP vs Anchor 10 11.86 18.64 5.00 0.019
LIME vs Anchor 10 5.87 18.64 2.00 0.006∗∗
LIME+SHAP vs Anchor 10 7.14 18.64 1.00 0.004∗∗
SHAP vs Anchor 10 8.85 18.64 0.00 0.002∗∗
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 7: Wilcoxon signed-rank test results comparing explainers (and their fused rule sets) to Anchor (Baseline) for $\bar{{CONF}}$.
Explainers set $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Anchor+LIME vs Anchor 10 0.85 0.84 17.00 0.32
Anchor+LIME+SHAP vs Anchor 10 0.87 0.84 9.00 0.06
Anchor+SHAP vs Anchor 10 0.88 0.84 4.00 0.014
LIME vs Anchor 10 0.81 0.84 13.00 0.16
LIME+SHAP vs Anchor 10 0.89 0.84 20.00 0.49
SHAP vs Anchor 10 0.86 0.84 26.00 0.92
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 8: Wilcoxon signed-rank test results comparing explainers (and their fused rule sets) to Anchor (Baseline) for $\bar{{COV}}$.
Explainers set $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Anchor+LIME vs Anchor 10 0.19 0.26 1.00 0.004∗∗
Anchor+LIME+SHAP vs Anchor 10 0.15 0.26 1.00 0.004∗∗
Anchor+SHAP vs Anchor 10 0.19 0.26 2.00 0.006∗∗
LIME vs Anchor 10 0.15 0.26 9.00 0.06
LIME+SHAP vs Anchor 10 0.12 0.26 6.00 0.027
SHAP vs Anchor 10 0.16 0.26 8.00 0.049
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 9: Wilcoxon signed-rank test results comparing explainers (and their fused rule sets) to Anchor (Baseline) for ER{{ER}}.
Explainers set $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Anchor+LIME vs Anchor 10 81.37 89.87 27.00 1.00
Anchor+LIME+SHAP vs Anchor 10 79.71 89.87 19.00 0.43
Anchor+SHAP vs Anchor 10 87.25 89.87 22.00 0.62
LIME vs Anchor 10 70.55 89.87 16.00 0.28
LIME+SHAP vs Anchor 10 81.04 89.87 24.00 0.77
SHAP vs Anchor 10 76.81 89.87 15.00 0.23
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 10: Wilcoxon signed-rank test results comparing explainers (and their fused rule sets) to Anchor (Baseline) for $\bar{F(n)}$.
Explainers set $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Anchor+LIME vs Anchor 10 7.34 1.55 0.00 0.002∗∗
Anchor+LIME+SHAP vs Anchor 10 9.89 1.55 1.00 0.004∗∗
Anchor+SHAP vs Anchor 10 4.56 1.55 1.00 0.004∗∗
LIME vs Anchor 10 9.12 1.55 1.00 0.004∗∗
LIME+SHAP vs Anchor 10 12.24 1.55 1.00 0.004∗∗
SHAP vs Anchor 10 4.18 1.55 2.00 0.006∗∗
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 11: Wilcoxon signed-rank test results comparing explainers (and their fused rule sets) to Anchor (Baseline) for Med(F(n)){Med}(F(n)).
Explainers set $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Anchor+LIME vs Anchor 10 3.29 1.75 0.00 0.002∗∗
Anchor+LIME+SHAP vs Anchor 10 5.96 1.75 1.00 0.004∗∗
Anchor+SHAP vs Anchor 10 2.92 1.75 1.00 0.004∗∗
LIME vs Anchor 10 4.50 1.75 4.00 0.014
LIME+SHAP vs Anchor 10 8.50 1.75 1.00 0.004∗∗
SHAP vs Anchor 10 3.88 1.75 5.50 0.027
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 12: Wilcoxon signed-rank test results comparing fusion methods (consisting of one or more explainers) to the Baseline explainer (Anchor, LIME or SHAP) for $\bar{M}$.
Methods $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Best vs Baseline 27 5.57 5.46 163.00 0.55
Intersection vs Baseline 27 3.92 5.46 166.00 0.59
Lasso vs Baseline 40 7.56 5.66 146.00 < 0.001∗∗∗
Lasso global vs Baseline 43 6.92 5.40 377.00 0.25
Union vs Baseline 27 1.76 5.46 48.00 < 0.001∗∗∗
Weighted vs Baseline 27 6.29 5.46 82.00 0.009∗∗
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 13: Wilcoxon signed-rank test results comparing fusion methods (consisting of one or more explainers) to the Baseline explainer (Anchor, LIME or SHAP) for $\bar{{CONF}}\cdot{ER}$.
Methods $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Best vs Baseline 27 95.59 70.97 0.00 < 0.001∗∗∗
Intersection vs Baseline 27 50.46 70.97 62.00 0.002∗∗
Lasso vs Baseline 40 78.49 73.31 114.00 < 0.001∗∗∗
Lasso global vs Baseline 43 28.73 70.97 141.00 < 0.001∗∗∗
Union vs Baseline 27 94.21 70.97 9.00 < 0.001∗∗∗
Weighted vs Baseline 27 76.89 70.97 79.00 0.007∗∗
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 14: Wilcoxon signed-rank test results comparing fusion methods (consisting of one or more explainers) to the Baseline explainer (Anchor, LIME or SHAP) for $\bar{{CONF}}\cdot\bar{{COV}}\cdot{ER}$.
Methods $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Best vs Baseline 27 10.00 11.26 152.00 0.39
Intersection vs Baseline 27 6.93 11.26 105.00 0.044
Lasso vs Baseline 40 10.38 11.02 245.00 0.026
Lasso global vs Baseline 43 9.85 10.02 459.00 0.87
Union vs Baseline 27 5.64 11.26 82.00 0.009∗∗
Weighted vs Baseline 28 12.15 11.31 95.00 0.013
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 15: Wilcoxon signed-rank test results comparing fusion methods (consisting of one or more explainers) to the Baseline explainer (Anchor, LIME or SHAP) for $\bar{{CONF}}$.
Methods $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Best vs Baseline 27 0.97 0.85 0.00 < 0.001∗∗∗
Intersection vs Baseline 27 0.83 0.85 85.00 0.011
Lasso vs Baseline 40 0.88 0.84 178.00 0.001∗∗
Lasso global vs Baseline 43 0.68 0.85 72.00 < 0.001∗∗∗
Union vs Baseline 27 0.94 0.85 47.00 < 0.001∗∗∗
Weighted vs Baseline 27 0.88 0.85 41.00 < 0.001∗∗∗
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 16: Wilcoxon signed-rank test results comparing fusion methods (consisting of one or more explainers) to the Baseline explainer (Anchor, LIME or SHAP) for $\bar{{COV}}$.
Methods $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Best vs Baseline 27 0.11 0.17 38.00 < 0.001∗∗∗
Intersection vs Baseline 27 0.16 0.17 178.00 0.80
Lasso vs Baseline 40 0.17 0.18 395.00 0.85
Lasso global vs Baseline 42 0.34 0.19 2.00 < 0.001∗∗∗
Union vs Baseline 27 0.06 0.17 16.00 < 0.001∗∗∗
Weighted vs Baseline 27 0.17 0.17 172.00 0.70
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 17: Wilcoxon signed-rank test results comparing fusion methods (consisting of one or more explainers) to the Baseline explainer (Anchor, LIME or SHAP) for ER{{ER}}.
Methods $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Best vs Baseline 21 100.00 86.67 0.00 < 0.001∗∗∗
Intersection vs Baseline 27 70.00 90.00 65.00 0.002∗∗
Lasso vs Baseline 25 93.02 90.00 26.00 < 0.001∗∗∗
Lasso global vs Baseline 43 53.43 91.67 241.00 0.004∗∗
Union vs Baseline 21 100.00 86.67 0.00 < 0.001∗∗∗
Weighted vs Baseline 26 89.60 90.39 148.00 0.50
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 18: Wilcoxon signed-rank test results comparing fusion methods (consisting of one or more explainers) to the Baseline explainer (Anchor, LIME or SHAP) for $\bar{F(n)}$.
Methods $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Best vs Baseline 27 27.50 23.59 90.00 0.016
Intersection vs Baseline 27 6.07 23.59 40.00 < 0.001∗∗∗
Lasso vs Baseline 40 5.05 10.65 19.00 < 0.001∗∗∗
Lasso global vs Baseline 43 1.02 9.33 46.00 < 0.001∗∗∗
Union vs Baseline 27 42.61 23.59 0.00 < 0.001∗∗∗
Weighted vs Baseline 27 23.90 23.59 128.00 0.15
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 19: Wilcoxon signed-rank test results comparing fusion methods (consisting of one or more explainers) to the Baseline explainer (Anchor, LIME or SHAP) for Med(F(n)){Med}(F(n)).
Methods $\mathbf{N}$ $\mathbf{Median_{1}}$ $\mathbf{Median_{2}}$ $\mathbf{W}$ $\mathbf{p}$
Best vs Baseline 26 26.50 21.58 61.50 0.003∗∗
Intersection vs Baseline 27 2.25 20.50 34.00 < 0.001∗∗∗
Lasso vs Baseline 28 4.83 21.58 44.00 < 0.001∗∗∗
Lasso global vs Baseline 42 0.67 7.17 73.00 < 0.001∗∗∗
Union vs Baseline 27 42.00 20.50 0.00 < 0.001∗∗∗
Weighted vs Baseline 26 17.33 21.58 132.50 0.28
*p(<0.05), **p(<0.01), ***p(<0.001)
A.1.4 Median Values of Metrics
Table 20: Summary of medians for fusion combinations across all metrics.
Fusion N $\bar{M}$ $\bar{{CONF}}\cdot{ER}$ $\bar{{CONF}}\cdot\bar{{COV}}\cdot{ER}$ $\bar{{CONF}}$ $\bar{{COV}}$ ER{{ER}} $\bar{F(n)}$ Med(F(n)){Med}(F(n))
Anchor 100 17.24 79.81 17.86 0.87 0.22 97.03 1.79 2.00
Anchor+LIME 60 10.62 84.73 12.31 0.90 0.17 95.92 6.66 2.75
Anchor+SHAP 60 10.44 90.47 12.01 0.92 0.15 97.69 4.91 2.25
Anchor+LIME+SHAP 60 6.92 84.52 8.55 0.91 0.12 93.59 8.49 4.75
SHAP 108 6.86 85.89 7.82 0.94 0.13 96.46 4.69 4.50
LIME+SHAP 162 5.51 86.03 7.47 0.94 0.10 92.91 12.54 8.50
LIME 112 5.05 64.79 6.05 0.87 0.13 85.90 10.90 5.38
Table 21: Summary of medians for fusion methods across all metrics.
Method N $\bar{M}$ $\bar{{CONF}}\cdot{ER}$ $\bar{{CONF}}\cdot\bar{{COV}}\cdot{ER}$ $\bar{{CONF}}$ $\bar{{COV}}$ ER{{ER}} $\bar{F(n)}$ Med(F(n)){Med}(F(n))
Lasso global 137 6.53 22.05 7.58 0.66 0.33 46.90 0.80 0.00
Weighted 137 6.15 82.19 11.33 0.90 0.15 93.83 22.74 13.00
Best 57 5.55 96.86 10.00 0.97 0.10 100.00 28.42 26.00
Lasso 137 5.28 91.05 9.35 0.94 0.12 100.00 6.69 6.00
Baseline 80 4.73 76.33 8.94 0.86 0.14 93.75 22.54 20.50
Intersection 57 3.92 51.32 6.29 0.81 0.13 70.00 5.76 1.25
Union 57 1.67 94.49 5.27 0.94 0.06 100.00 42.61 42.00
A.2 Additional Figures
A.2.1 Plots for Baseline (single) Explainers
(a) $\bar{M}$ (Objective function)
(b) $\overline{\text{CONF}}\cdot ER$ (Avg. Conf.*Expl. R.)
(c) $\overline{\text{CONF}}$ (Average Confidence)
(d) $\overline{\text{COV}}$ (Average Coverage)
(e) ERER (Explained Ratio)
(f) $\overline{F(n)}$ (Average Feature Count)
Figure 8: Performance comparison of baseline (single) explainers across key metrics. The metrics displayed are: $\bar{M}$ (Objective function), $\overline{\text{CONF}}\cdot ER$ (Average Confidence * Explained Ratio), $\overline{\text{CONF}}$ (Average Confidence), $\overline{\text{COV}}$ (Average Coverage), ERER (Explained Ratio), and $\overline{F(n)}$ (Average Feature Count). Each plot visualizes the metric distribution for Anchor, LIME, and SHAP under the baseline condition.
A.2.2 Comparison of Baselines and Fusion Explainers
(a) Objective function performance ($\bar{M}$).
(b) Combined metric $\overline{\text{CONF}}\cdot ER$ (Confidence $\cdot$ Explained Ratio).
(c) Metric $\overline{\text{COV}}$ (Coverage).
(d) Average feature count ($\bar{F(n)}$).
Figure 9: Single-explainer baselines (orange) vs fusion methods (blue). Boxplots are annotated with mean, standard deviation, and median. Circles denote outliers.
Figure 10: Objective function performance ($\bar{M}$): Comparison of single-explainer baselines (orange) and fusion methods (blue). Circles denote outliers.
Figure 11: Combined metric $\overline{\text{CONF}}\cdot ER$ (Confidence $\cdot$ Explained Ratio): Comparison of single-explainer baselines (orange) and fusion methods (blue). Circles denote outliers.
Figure 12: Metric $\overline{\text{COV}}$ (Coverage): Comparison of coverage metrics for single-explainer baselines (orange) and fusion methods (blue). Circles denote outliers.
Figure 13: Average feature count ($\bar{F(n)}$): Distribution of feature counts used by single-explainer baselines (orange) and fusion methods (blue). Circles denote outliers.
A.2.3 Additional Critical Difference Diagrams
(a) Critical Difference Diagram for $\bar{M}$ (Final Metric).
(b) Critical Difference Diagram for ERER (Explained Ratio).
(c) Critical Difference Diagram for $F(\bar{n})$ (Average Feature Count).
Figure 14: Critical Difference Diagrams for different fusion methods and baselines across metrics. Subfigure (a) shows $\bar{M}$ (Final Metric), (b) illustrates ERER (Explained Ratio), and (c) presents $F(\bar{n})$ (Average Feature Count).
A.3 How to read the rules.
Each panel shows the raw time series. Red highlights mark time points that satisfy a rule. Dashed vertical lines indicate the start and end of an interval $\tau_{1},\tau_{2}\,]$.
For univariate data, a rule is a finite conjunction of point-wise constraints on selected indices $t_{j_{k}}$, each constraining the raw value to a half-open interval: $x_{t_{j_{k}}}\in(\ell_{k},u_{k}]$. For multivariate data, we indicate the feature/channel by a superscript: $x^{(c_{j_{k}})}_{t_{j_{k}}}\in(\ell_{k},u_{k}]$. Intervals are obtained by our numeric-to-rules procedure (see Section 3.3).
$R_{i}:\quad\left\{\begin{aligned} &x_{t_{j_{1}}}\in(\ell_{1},u_{1}]\\ &x_{t_{j_{2}}}\in(\ell_{2},u_{2}]\\ &\;\vdots\\ &x_{t_{j_{m}}}\in(\ell_{m},u_{m}]\end{aligned}\right.\;\Rightarrow\;\begin{aligned} \hat{y}&=\alpha,\\ \mathrm{CONF}(R_{i})&=\gamma,\\ \mathrm{COV}(R_{i})&=\kappa.\end{aligned}$
We report confidence (precision) and coverage for each rule. Rules are human-readable proxies of model behaviour: they show where and what the classifier relied on, but they are not causal claims. Prototypes and class averages, when shown, give shape context. The figures are actionable: a practitioner can verify the inequality on the highlighted segment or inspect the waveform there.
A.4 Dataset-level exemplar visualisations
A.4.1 Dataset-level exemplar visualisations for ECG200
Visualizations intuitively link numeric feature attribution explanations to TS structure, aiding interpretability. Our framework transforms numeric explanations into rules consisting of features and intervals of their applicability, which can be visualized on TS plots and aid revealing patterns and decision-making.
(a) ECG200 with Anchor (baseline)
(b) ECG200 with LIME (baseline)
(c) ECG200 with SHAP (baseline)
(d) ECG200 with Anchor+LIME+SHAP (Lasso fusion)
Figure 15: Visual explanations for univariate ECG200 data (two classes).
Figure 15 shows LIME and SHAP explanations converted into rule-based intervals for ECG200 dataset (univariate, two classes), as well as Anchor (native rule-based). Each subplot consists of five plots per row, where each row represents visualizations for one class. The next row contains the second class. The first three columns display green-plotted instances that are maximally diverse, while the fourth column represents the most average, prototypical observation (orange-plotted). The last (green) column shows the class average. Red vertical lines mark feature intervals where time points satisfy rules. LIME targets mainly the initial waveform sections, whereas SHAP highlights central intervals. Anchor+LIME+SHAP Lasso fusion merges both methods, improving coverage and confidence ratio. This slightly increases the median feature count from 4 to 5. Notably, Observation2 in Class1, under-unexplained by LIME, is covered by fusion. Despite added complexity, fusion remains computationally efficient for TS.
A.4.2 Dataset-level exemplar visualisations for Plane
The Plane dataset (univariate, 6 classes) highlights differences in rule intervals from Anchor, SHAP, LIME, and their fusion (Figures 16-19, Appendix). Anchor uses fewer variables, but its red-marked intervals are often too wide. This reduces informativeness, as broad intervals fail to capture nuances of TS shape. SHAP and LIME use more variables, while their interval method (Section 3.3) generates narrower ranges via multidimensional perturbations. Narrower intervals may improve interpretability by focusing attention on specific waveform segments. However, overly narrow intervals risk overfitting. Our fusion step trades off precision, coverage, and simplicity to mitigate this. Fusing Anchor, SHAP, and LIME with "best" method matches SHAP’s baseline performance with significantly fewer variables.
Figure 16: Plane dataset with Anchor (baseline).
Figure 17: Plane dataset with SHAP (baseline).
Figure 18: Plane dataset with LIME (baseline).
Figure 19: Plane dataset with Anchor+LIME+SHAP (Best Combination).
B Supplementary Materials
Friedman and Nemenyi: To compare multiple methods statistically, we applied the Friedman test, followed by post-hoc Nemenyi comparisons. The Friedman test results provide insights into the statistical significance of differences across methods. Table 22 reports results at the method level, considering different fusion strategies (e.g., Baseline, Best, Intersection, Weighted). Table 23 presents results for all rule sets, accounting for the combinations of explainers they comprise, both single and compound. For pairwise comparisons, we employed the Nemenyi test. The first set of tables, from Table 24 to Table 31, evaluates fusion combinations in terms of explainer types. The second set, from Table 32 to Table 39, focuses on comparisons across fusion methods.
B.1 Friedman Tests
Table 22: Friedman test results for all fusion methods (e.g. Baseline, Best, Intersection, … , Weighted) across metrics.
Metric $\chi^{2}$ $p$
$\bar{M}$ 41.46 < 0.001∗∗∗
$\bar{{CONF}}\cdot{ER}$ 130.10 < 0.001∗∗∗
$\bar{{CONF}}\cdot\bar{{COV}}\cdot{ER}$ 29.49 < 0.001∗∗∗
$\bar{{CONF}}$ 95.69 < 0.001∗∗∗
$\bar{{COV}}$ 110.02 < 0.001∗∗∗
ER{{ER}} 126.00 < 0.001∗∗∗
$\bar{F(n)}$ 136.22 < 0.001∗∗∗
Med(F(n)){Med}(F(n)) 128.66 < 0.001∗∗∗
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 23: Friedman test results for all fused rule sets (taking into account explainers combinations they are composed of, e.g. Anchor, Anchor+LIME, Anchor+LIME+SHAP, etc.) across metrics.
Metric $\chi^{2}$ $p$
$\bar{M}$ 36.09 < 0.001∗∗∗
$\bar{{CONF}}\cdot{ER}$ 17.31 0.008∗∗
$\bar{{CONF}}\cdot\bar{{COV}}\cdot{ER}$ 30.64 < 0.001∗∗∗
$\bar{{CONF}}$ 19.16 0.004∗∗
$\bar{{COV}}$ 13.50 0.036
ER{{ER}} 13.41 0.037
$\bar{F(n)}$ 27.26 < 0.001∗∗∗
Med(F(n)){Med}(F(n)) 28.03 < 0.001∗∗∗
*p(<0.05), **p(<0.01), ***p(<0.001)
B.2 Nemenyi Tests
Table 24: Nemenyi test results for fusion combinations for $\bar{M}$. P-values for pairwise comparisons.
Rules Combination Anchor Anchor+LIME Anchor+LIME+SHAP Anchor+SHAP LIME LIME+SHAP SHAP
Anchor 1.00 0.25 0.023* 0.69 0.001** 0.001** 0.002**
Anchor+LIME 0.25 1.00 0.90 0.90 0.13 0.31 0.63
Anchor+LIME+SHAP 0.023* 0.90 1.00 0.63 0.63 0.87 0.90
Anchor+SHAP 0.69 0.90 0.63 1.00 0.016* 0.06 0.21
LIME 0.001** 0.13 0.63 0.016* 1.00 0.90 0.90
LIME+SHAP 0.001** 0.31 0.87 0.06 0.90 1.00 0.90
SHAP 0.002** 0.63 0.90 0.21 0.90 0.90 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 25: Nemenyi test results for fusion combinations for $\bar{{CONF}}\cdot{ER}$. P-values for pairwise comparisons.
Rules Combination Anchor Anchor+LIME Anchor+LIME+SHAP Anchor+SHAP LIME LIME+SHAP SHAP
Anchor 1.00 0.90 0.90 0.69 0.37 0.90 0.90
Anchor+LIME 0.90 1.00 0.56 0.90 0.043* 0.90 0.69
Anchor+LIME+SHAP 0.90 0.56 1.00 0.21 0.87 0.90 0.90
Anchor+SHAP 0.69 0.90 0.21 1.00 0.005** 0.87 0.31
LIME 0.37 0.043* 0.87 0.005** 1.00 0.21 0.75
LIME+SHAP 0.90 0.90 0.90 0.87 0.21 1.00 0.90
SHAP 0.90 0.69 0.90 0.31 0.75 0.90 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 26: Nemenyi test results for fusion combinations for $\bar{{CONF}}\cdot\bar{{COV}}\cdot{ER}$. P-values for pairwise comparisons.
Rules Combination Anchor Anchor+LIME Anchor+LIME+SHAP Anchor+SHAP LIME LIME+SHAP SHAP
Anchor 1.00 0.87 0.06 0.75 0.004** 0.001** 0.001**
Anchor+LIME 0.87 1.00 0.63 0.90 0.17 0.08 0.08
Anchor+LIME+SHAP 0.06 0.63 1.00 0.75 0.90 0.90 0.90
Anchor+SHAP 0.75 0.90 0.75 1.00 0.25 0.13 0.13
LIME 0.004** 0.17 0.90 0.25 1.00 0.90 0.90
LIME+SHAP 0.001** 0.08 0.90 0.13 0.90 1.00 0.90
SHAP 0.001** 0.08 0.90 0.13 0.90 0.90 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 27: Nemenyi test results for fusion combinations for $\bar{{CONF}}$. P-values for pairwise comparisons.
Rules Combination Anchor Anchor+LIME Anchor+LIME+SHAP Anchor+SHAP LIME LIME+SHAP SHAP
Anchor 1.00 0.90 0.90 0.50 0.31 0.90 0.90
Anchor+LIME 0.90 1.00 0.90 0.90 0.043* 0.90 0.90
Anchor+LIME+SHAP 0.90 0.90 1.00 0.90 0.043* 0.90 0.90
Anchor+SHAP 0.50 0.90 0.90 1.00 0.001** 0.90 0.50
LIME 0.31 0.043* 0.043* 0.001** 1.00 0.043* 0.31
LIME+SHAP 0.90 0.90 0.90 0.90 0.043* 1.00 0.90
SHAP 0.90 0.90 0.90 0.50 0.31 0.90 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 28: Nemenyi test results for fusion combinations for $\bar{{COV}}$. P-values for pairwise comparisons.
Rules Combination Anchor Anchor+LIME Anchor+LIME+SHAP Anchor+SHAP LIME LIME+SHAP SHAP
Anchor 1.00 0.44 0.13 0.31 0.17 0.011* 0.13
Anchor+LIME 0.44 1.00 0.90 0.90 0.90 0.75 0.90
Anchor+LIME+SHAP 0.13 0.90 1.00 0.90 0.90 0.90 0.90
Anchor+SHAP 0.31 0.90 0.90 1.00 0.90 0.87 0.90
LIME 0.17 0.90 0.90 0.90 1.00 0.90 0.90
LIME+SHAP 0.011* 0.75 0.90 0.87 0.90 1.00 0.90
SHAP 0.13 0.90 0.90 0.90 0.90 0.90 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 29: Nemenyi test results for fusion combinations for ER{{ER}}. P-values for pairwise comparisons.
Rules Combination Anchor Anchor+LIME Anchor+LIME+SHAP Anchor+SHAP LIME LIME+SHAP SHAP
Anchor 1.00 0.90 0.50 0.90 0.31 0.90 0.63
Anchor+LIME 0.90 1.00 0.44 0.90 0.25 0.90 0.56
Anchor+LIME+SHAP 0.50 0.44 1.00 0.25 0.90 0.87 0.90
Anchor+SHAP 0.90 0.90 0.25 1.00 0.13 0.90 0.37
LIME 0.31 0.25 0.90 0.13 1.00 0.69 0.90
LIME+SHAP 0.90 0.90 0.87 0.90 0.69 1.00 0.90
SHAP 0.63 0.56 0.90 0.37 0.90 0.90 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 30: Nemenyi test results for fusion combinations for $\bar{F(n)}$. P-values for pairwise comparisons.
Rules Combination Anchor Anchor+LIME Anchor+LIME+SHAP Anchor+SHAP LIME LIME+SHAP SHAP
Anchor 1.00 0.023* 0.016* 0.56 0.016* 0.001** 0.50
Anchor+LIME 0.023* 1.00 0.90 0.75 0.90 0.75 0.81
Anchor+LIME+SHAP 0.016* 0.90 1.00 0.69 0.90 0.81 0.75
Anchor+SHAP 0.56 0.75 0.69 1.00 0.69 0.06 0.90
LIME 0.016* 0.90 0.90 0.69 1.00 0.81 0.75
LIME+SHAP 0.001** 0.75 0.81 0.06 0.81 1.00 0.08
SHAP 0.50 0.81 0.75 0.90 0.75 0.08 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 31: Nemenyi test results for fusion combinations for Med(F(n)){Med}(F(n)). P-values for pairwise comparisons.
Rules Combination Anchor Anchor+LIME Anchor+LIME+SHAP Anchor+SHAP LIME LIME+SHAP SHAP
Anchor 1.00 0.31 0.011* 0.78 0.050* 0.001** 0.44
Anchor+LIME 0.31 1.00 0.87 0.90 0.90 0.13 0.90
Anchor+LIME+SHAP 0.011* 0.87 1.00 0.40 0.90 0.81 0.75
Anchor+SHAP 0.78 0.90 0.40 1.00 0.69 0.014* 0.90
LIME 0.050* 0.90 0.90 0.69 1.00 0.53 0.90
LIME+SHAP 0.001** 0.13 0.81 0.014* 0.53 1.00 0.08
SHAP 0.44 0.90 0.75 0.90 0.90 0.08 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 32: Nemenyi test results for fusion methods for $\bar{M}$. P-values for pairwise comparisons.
Method Baseline Best Intersection Lasso Lasso global Union Weighted
Baseline 1.00 0.015* 0.56 0.49 0.56 0.90 0.90
Best 0.015* 1.00 0.001** 0.001** 0.001** 0.001** 0.018*
Intersection 0.56 0.001** 1.00 0.90 0.90 0.90 0.53
Lasso 0.49 0.001** 0.90 1.00 0.90 0.90 0.45
Lasso global 0.56 0.001** 0.90 0.90 1.00 0.90 0.53
Union 0.90 0.001** 0.90 0.90 0.90 1.00 0.90
Weighted 0.90 0.018* 0.53 0.45 0.53 0.90 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 33: Nemenyi test results for fusion methods for $\bar{{CONF}}\cdot{ER}$. P-values for pairwise comparisons.
Method Baseline Best Intersection Lasso Lasso global Union Weighted
Baseline 1.00 0.001** 0.21 0.001** 0.47 0.001** 0.80
Best 0.001** 1.00 0.003** 0.33 0.001** 0.90 0.001**
Intersection 0.21 0.003** 1.00 0.64 0.001** 0.001** 0.90
Lasso 0.001** 0.33 0.64 1.00 0.001** 0.020* 0.11
Lasso global 0.47 0.001** 0.001** 0.001** 1.00 0.001** 0.018*
Union 0.001** 0.90 0.001** 0.020* 0.001** 1.00 0.001**
Weighted 0.80 0.001** 0.90 0.11 0.018* 0.001** 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 34: Nemenyi test results for fusion methods for $\bar{{CONF}}\cdot\bar{{COV}}\cdot{ER}$. P-values for pairwise comparisons.
Method Baseline Best Intersection Lasso Lasso global Union Weighted
Baseline 1.00 0.90 0.008** 0.18 0.37 0.10 0.64
Best 0.90 1.00 0.001** 0.013* 0.044* 0.006** 0.14
Intersection 0.008** 0.001** 1.00 0.90 0.75 0.90 0.49
Lasso 0.18 0.013* 0.90 1.00 0.90 0.90 0.90
Lasso global 0.37 0.044* 0.75 0.90 1.00 0.90 0.90
Union 0.10 0.006** 0.90 0.90 0.90 1.00 0.90
Weighted 0.64 0.14 0.49 0.90 0.90 0.90 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 35: Nemenyi test results for fusion methods for $\bar{{CONF}}$. P-values for pairwise comparisons.
Method Baseline Best Intersection Lasso Lasso global Union Weighted
Baseline 1.00 0.001** 0.14 0.016* 0.90 0.001** 0.90
Best 0.001** 1.00 0.14 0.54 0.001** 0.77 0.001**
Intersection 0.14 0.14 1.00 0.90 0.004** 0.001** 0.35
Lasso 0.016* 0.54 0.90 1.00 0.001** 0.022* 0.06
Lasso global 0.90 0.001** 0.004** 0.001** 1.00 0.001** 0.65
Union 0.001** 0.77 0.001** 0.022* 0.001** 1.00 0.001**
Weighted 0.90 0.001** 0.35 0.06 0.65 0.001** 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 36: Nemenyi test results for fusion methods for $\bar{{COV}}$. P-values for pairwise comparisons.
Method Baseline Best Intersection Lasso Lasso global Union Weighted
Baseline 1.00 0.001** 0.65 0.90 0.001** 0.23 0.90
Best 0.001** 1.00 0.001** 0.001** 0.001** 0.35 0.001**
Intersection 0.65 0.001** 1.00 0.75 0.008** 0.002** 0.90
Lasso 0.90 0.001** 0.75 1.00 0.001** 0.16 0.90
Lasso global 0.001** 0.001** 0.008** 0.001** 1.00 0.001** 0.001**
Union 0.23 0.35 0.002** 0.16 0.001** 1.00 0.016*
Weighted 0.90 0.001** 0.90 0.90 0.001** 0.016* 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 37: Nemenyi test results for fusion methods for ER{{ER}}. P-values for pairwise comparisons.
Method Baseline Best Intersection Lasso Lasso global Union Weighted
Baseline 1.00 0.001** 0.49 0.001** 0.69 0.001** 0.39
Best 0.001** 1.00 0.001** 0.13 0.001** 0.90 0.001**
Intersection 0.49 0.001** 1.00 0.29 0.011* 0.001** 0.90
Lasso 0.001** 0.13 0.29 1.00 0.001** 0.13 0.39
Lasso global 0.69 0.001** 0.011* 0.001** 1.00 0.001** 0.006**
Union 0.001** 0.90 0.001** 0.13 0.001** 1.00 0.001**
Weighted 0.39 0.001** 0.90 0.39 0.006** 0.001** 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 38: Nemenyi test results for fusion methods for $\bar{F(n)}$. P-values for pairwise comparisons.
Method Baseline Best Intersection Lasso Lasso global Union Weighted
Baseline 1.00 0.001** 0.10 0.90 0.10 0.001** 0.027*
Best 0.001** 1.00 0.001** 0.001** 0.001** 0.20 0.001**
Intersection 0.10 0.001** 1.00 0.11 0.001** 0.29 0.90
Lasso 0.90 0.001** 0.11 1.00 0.08 0.001** 0.033*
Lasso global 0.10 0.001** 0.001** 0.08 1.00 0.001** 0.001**
Union 0.001** 0.20 0.29 0.001** 0.001** 1.00 0.56
Weighted 0.027* 0.001** 0.90 0.033* 0.001** 0.56 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)
Table 39: Nemenyi test results for fusion methods for Med(F(n)){Med}(F(n)). P-values for pairwise comparisons.
Method Baseline Best Intersection Lasso Lasso global Union Weighted
Baseline 1.00 0.001** 0.05 0.49 0.43 0.001** 0.005**
Best 0.001** 1.00 0.001** 0.001** 0.001** 0.20 0.001**
Intersection 0.05 0.001** 1.00 0.90 0.001** 0.18 0.90
Lasso 0.49 0.001** 0.90 1.00 0.002** 0.009** 0.58
Lasso global 0.43 0.001** 0.001** 0.002** 1.00 0.001** 0.001**
Union 0.001** 0.20 0.18 0.009** 0.001** 1.00 0.56
Weighted 0.005** 0.001** 0.90 0.58 0.001** 0.56 1.00
*p(<0.05), **p(<0.01), ***p(<0.001)